### v1

# Demystifying Domain-adaptive Post-training for Financial LLMs

Abstract
Domain-adaptive post-training of large language models (LLMs) has emerged as a promising approach for specialized domains such as medicine and finance. However, significant challenges remain in identifying optimal adaptation criteria and training strategies across varying data and model configurations. To address these challenges, we introduce FinDaP, a systematic and fine-grained investigation into domain-adaptive post-training of LLMs for the finance domain. Our approach begins by identifying the core capabilities required for the target domain and designing a comprehensive evaluation suite aligned with these needs. We then analyze the effectiveness of key post-training stages, including continual pre-training, instruction tuning, and preference alignment. Building on these insights, we propose an effective training recipe centered on a novel preference data distillation method, which leverages process signals from a generative reward model. The resulting model, Llama-Fin, achieves state-of-the-art performance across a wide range of financial tasks. Our analysis also highlights how each post-training stage contributes to distinct capabilities, uncovering specific challenges and effective solutions, providing valuable insights for domain adaptation of LLMs.
Project page: https://github.com/SalesforceAIResearch/FinDap
1 Introduction
While large language models (LLMs) demonstrate strong generalization across diverse tasks, they often struggle in specialized domains such as finance, law, and healthcare. Consequently, domain-adaptive post-training of LLMs has garnered significant attention recently (Colombo et al., 2024b; Xie et al., 2024a). In the earlier days of language models, continual pre-training or CPT was the dominant strategy. This involved further training a pre-trained model on domain-specific plain text and then fine-tuning it for individual tasks (Gururangan et al., 2020; Ke et al., 2023). While prompt engineering with zero- or few-shot examples has emerged as a convenient way to adapt general-purpose LLMs to new tasks, recent methods focus on fine-tuning model weights to transform them into domain experts (Colombo et al., 2024a; Chen et al., 2023a; Li et al., 2023). [^1] This shift is understandable not only because training typically leads to better results on specialized tasks, but also because training has become more affordable due to the growing number of techniques that enable efficient training. For example, the training cost was over \$4M for GPT-3 (davinci), but it has recently dropped to around \$0.8M to train a GPT-3 on-par model like Phi3.5  (Maslej et al., 2024).
Today, domain-adaptive post-training has evolved into a collection of techniques including CPT, instruction-tunning (IT) and preference alignment (PA). Building on this trend, this work focuses on adapting LLMs to specific domains through parameter training, which complements semi-parametric methods that leverage external knowledge, such as in retrieval-augmented generation or RAG (Lewis et al., 2020; Ke et al., 2024). Our focus is also different from general post-training, as the goal is not to develop another general-purpose frontier LLM but to create specialized, expert-level LLMs tailored to a specific domain.
Despite the need and aforementioned advancements in domain-specific post-training, there is still no systematic study on what makes a good domain-specific LLM (see Table 2). In this paper, we wish to address the following research questions:
Research Questions Given a pre-trained LLM with strong general capabilities (e.g., Llama3-8b-instruct), how to effectively adapt it to a target domain by post-training? What criteria are desirable for successful adaptation? What are effective training recipes with respect to data and model?
A common criterion for evaluating domain-specific LLMs is to test their performance on a set of domain-specific tasks (Bhatia et al., 2024; Xie et al., 2024b). However, this simplified approach can often misalign with our broader expectations for a domain-expert LLM. The absence of clear and comprehensive evaluation criteria has significantly hindered the development of robust training recipes, encompassing both data and model strategies. Furthermore, it remains uncertain whether simply using a larger LLM will yield better performance in a specific domain compared to a carefully post-trained, domain-adaptive LLM.
To address these questions methodically, we introduce FinDaP, a novel finance-specific post-training framework comprising a comprehensive evaluation framework, state-of-the-art model checkpoints and a training recipe. We use the finance domain as a case study to perform domain-adaptive post-training on an instruction-tuned LLM, Llama3-8b-instruct (LLaMA, 2024). Figure 1 outlines our approach. We begin by identifying a set of core capabilities that a domain-expert LLM should possess such as knowledge about concepts and tasks in the target domain and the associated reasoning processes (§2.1). This identification is critical as it informs the entire exploration process. Based on these expected capabilities, we develop an evaluation framework (§2.2) to establish clear performance goals and guide model improvement over a selection of development and unseen tasks.
The training recipe involves two stages. In the first stage (§3.2 & §3.3), we conduct rigorous and principled experiments to determine the optimal data and model recipes for CPT and IT, resulting in the ‘CPT+IT’ checkpoint. Subsequently, we further train the model using curated on-policy preference data for PA. We propose a novel approach to construct preference data from the process signals given by a generative reward model. Our best performing recipe yields Llama-Fin that outperforms all considered baselines, including large open models at the 70B scale and proprietary models like GPT-4o, on unseen tasks that are similar to the training data. Even on unseen and novel tasks never encountered in training, Llama-Fin remains competitive and consistently outperforms its base model across all identified capabilities. In summary, our key contributions are:
  • A comprehensive guidance for domain-specific post-training, including identification of capabilities, evaluation, data and model recipe design.
  • A systematic exploration on each stage of domain-specific post-training, with an emphasis on the goals, challenges and effective approaches at each stage.
  • A novel preference alignment approach that constructs preference data using on-policy trajectories guided by outcome and process signals from a generative reward model.
  • A new state-of-the-art finance LLM (Llama-Fin) at the 8b parameter scale.
  • Extensive evaluation with an open-source leaderboard, checkpoints, data and model recipes.
Figure 1: An overview of FinDaP. Left: we first identify the core expected capabilities for the target domain and then curate texts and prompts for training and evaluation. Right: the top shows our training strategies. For each training stage, we use development set to select the best model. After training, we use unseen set to demonstrate the effectiveness of Llama-Fin.
2 FinDaP Setup
In FinDaP, we first identify the desired capabilities and design the corresponding evaluation measures, and then formulate the training method.
2.1 FinDaP Capabilities
We begin by illustrating the capabilities that are desirable for a domain-specific LLM. Specifically, we focus on the following core capabilities:
$\bullet$ Domain specific concepts. A domain typically includes its own specific concepts. For example, ‘bond’ in finance refers to a loan agreement between an investor and a borrower. Adapting the LLM to domain-specific concepts is crucial, as these concepts form the fundamental building blocks of domain knowledge. However, this adaptation should not come at the cost of losing knowledge about general concepts, which are essential for both domain-specific and general tasks.
$\bullet$ Domain specific tasks. While many NLP tasks, such as NER or sentiment analysis, are shared across different domains, a domain typically has its own tasks. For example, stock movement detection is primarily found in finance. Adapting LLMs to these domain-specific tasks is important, as it demonstrates how they can leverage domain-specific concepts to solve tailored tasks effectively.
$\bullet$ Reasoning. For complex tasks, reasoning with concepts is a highly desired capability in LLMs. For example, in finance, the LLM is often required to analyze a company’s financial report, involving extensive reasoning, particularly mathematical reasoning, to compute key financial concepts such as market rate or earnings per share.
$\bullet$ Instruction-Following (IF) and chat. This is a core capability for both general and domain-specific LLMs, as tasks are often presented in the form of instruction following or conversation.
$\bullet$ Others. Additionally, domains may vary significantly in their sensitivity. For instance, the medical domain is highly sensitive, requiring utmost accuracy and strict adherence to ethical considerations. In contrast, domains such as entertainment may have more relaxed requirements. Another important consideration is multi-modality, as some domains require handling multiple types of input and output formats. For example, the healthcare domain may involve processing medical images alongside textual reports, while the e-commerce domain may integrate product descriptions, images, and customer reviews into a unified response. Similarly, scientific research often combines charts, graphs, and textual analysis to present findings effectively. While we acknowledge these additional aspects, we leave those for future work and concentrate on the four primary capabilities discussed above.
2.2 FinDaP Evaluation
With the above breakdown of capabilities, our evaluation framework consists of a suite for assessing these capabilities using development sets and unseen (held-out) evaluation sets. Our development set is directly split from the training data at each stage. Table 1 outlines the capabilities and the evaluation benchmarks selected to cover these capabilities. Crucially, we did not examine scores on our unseen set while developing the models, which allows us to observe how much we may have overfitted to particular evaluations in our decisions around training recipe. For the unseen tasks (Table 1), we manually review each individual dataset and have the following considerations.
Capability Domain Task Evaluation Dataset Size Reference
Unseen - Similar
Tasks Finance Sentiment Analysis FPB 970 Malo et al. (2014)
FiQA SA 235 Maia et al. (2018)
Monetary policy Stance FOMC 496 Shah et al. (2023)
Named entity recognition NER 98 Alvarado et al. (2015)
Abstractive Summarization EDTSUM 2,000 Zhou et al. (2021)
Total 3,799
Unseen - Novel
Concept General Knowledge Recall MMLU 14,042 (Hendrycks et al., 2021)
AI2-ARC 3,548 Clark et al. (2018)
Nq-open 7,842 Kwiatkowski et al. (2019)
Finance MMLU-Finance 1,460 -
Tasks Finance Extractive Summarization Flare-ECTSUM 495 Mukherjee et al. (2022)
ESG Issue Classification MLESG 300 Chen et al. (2023b)
Rumour Detection MA 500 Yang et al. (2020)
Stock Movement Prediction SM-Bigdata 1,470 Soun et al. (2022)
SM-ACL 3,720 Xu and Cohen (2018)
SM-CIKM 1,140 Wu et al. (2018)
Fraud Detection CRA-CCF 2,280 Feng et al. (2024)
CRA-CCFraud 2,100 Feng et al. (2024)
Credit Scoring Flare-German 200 Hofmann (1994)
Flare-Astralian 139 Quinlan (1987)
CRA-LendingClub 2,690 Feng et al. (2024)
Distress Identification CRA-Polish 1,740 Feng et al. (2024)
CRA-Taiwan 1,370 Feng et al. (2024)
Claim Analysis CRA-ProroSeguro 2,380 Feng et al. (2024)
CRA-TravelInsurance 2,530 Feng et al. (2024)
Tabular QA Flare-TATQA 1,670 Zhu et al. (2021)
Open QA Finance Bench 150 Islam et al. (2023)
IF/Chat General Precise IF MT-bench 80 Zheng et al. (2023)
Reasoning Math Reasoning MathQA 2,985 Amini et al. (2019a)
General Social Reasoning Social-IQA 2,636 Welbl et al. (2017)
Common Reasoning Open-book-qa 500 Mihaylov et al. (2018)
Hellaswag 10,003 Zellers et al. (2019)
Winogrande 1,767 Sakaguchi et al. (2019)
PIQA 3,000 Bisk et al. (2020)
Finance Exam CFA-Easy 1,030 Link
CFA-Challenge 90 -
Total 91,872
Table 1: Summary of our evaluation dataset. New datasets released with FinDaP are color-highlighted for emphasis.
$\bullet$ Benchmarking tasks. Corresponding to the capabilities, we consider a diverse set of benchmarking tasks. For concepts, we include knowledge tasks in the general domain, such as AI2-ARC (Clark et al., 2018), as well as in finance, such as MMLU-Finance (Hendrycks et al., 2021). For tasks, we consider general tasks, such as Social-IQA (Welbl et al., 2017), and domain-specific tasks, such as MLESG (Chen et al., 2023b). Notably, we intentionally include a few financial tasks such as Flare-TATQA (Zhu et al., 2021) and SM-Bigdata (Soun et al., 2022) that require understanding of tabular data, as this data format is common in this domain. For IF/Chat capabilities, we utilize popular instruction-following benchmarks, such as MT-Bench (Zheng et al., 2023). For reasoning, we include general reasoning tasks, such as MathQA (Amini et al., 2019a) and Hellaswag common sense reasoning (Zellers et al., 2019), as well as domain-specific reasoning tasks, such as CRA-ProroSeguro claim analysis (Feng et al., 2024). We also construct a new benchmark on CFA-Challenge based on CFA Level III, one of the most challenging financial exams that requires comprehensive reasoning (Khamnuansin et al., 2024; Callanan et al., 2024).
$\bullet$ Evaluation method. We split our evaluation set into two types based on their exposure to Instruction tuning (IT) data (Table 1). The first type, Similar, includes tasks whose types have been encountered during training, even if the specific tasks themselves are unseen (e.g., a new NER task). The second type, Novel, includes tasks whose types have not been seen during training, representing entirely new challenges for the model (e.g., stock movement prediction). We use two different evaluation methods based on the nature of the benchmarks. For knowledge and NLP tasks (e.g., NER), we employ a straightforward direct answer evaluation. For reasoning tasks (e.g., CFA-Challenge), we use a 0-shot chain-of-thought (CoT) (Wei et al., 2023) answer evaluation to enhance the reliability of our evaluation. This also exposes the reasoning path, allowing us to investigate the causes of incorrect answers and enabling a more fine-grained comparison across different models.
2.3 Comparison with Existing Finance LLMs
Table 2 compares Llama-Fin with popular open finance-specific LLMs: FinTral (Bhatia et al., 2024), PIXIU (Xie et al., 2023), FinLLM (Xie et al., 2024b), and AdaptLLM (Cheng et al., 2024). Among these models, only FinTral incorporates all three training stages: CPT, IT, and PA. However, unlike Llama-Fin, FinTral and other models lack a systematic design and fail to include carefully crafted components, as illustrated in Figure 1. For instance, none of these models have an in-depth analysis of design choices for different post-training stages, let alone introduce a novel approach for constructing preference data.
Finance LLM Capability Training Evaluation
Continue Pre-training Instruction Tuning Preference Alignment Task Method
General Domain Target Domain General Task Domain Task On-policy General Task Domain Task Reasoning Task Direct Answer CoT Answer
ORM PRM
FinTral Concept, Task
PIXIU Task
FinLLM Concept, Task
AdaptLLM Concept
Llama-Fin
Concept, IF/Chat,
Task, Reasoning
Table 2: Comparison between Llama-Fin with other open finance LLMs.
3 Model Training
As shown in Figure 1, FinDaP’s training starts with curated text and prompts (§3.1), followed by joint Continual Pre-training and Instruction Tuning (§3.4), and then Preference Alignment (§3.5).
3.1 Data
Text Curation. To introduce domain concepts while preserving general concepts, we curate texts for CPT. Table 3 summarizes the texts curation datasets. Specially, for general concepts, research has shown that a ‘small’ amount of general text (as little as 1%) can effectively mitigate the forgetting issue (Scialom et al., 2022). Therefore, we focus on collecting a relatively small but high-quality set of general-domain text. To achieve this, we use verifiable text, which is text written by humans and previously used in supervised tasks in the literature. Note that this contrasts with using unverifiable web text such as C4 (Raffel et al., 2020).
Capability Domain CPT Dataset Size Reference
Concept General NaturalInstrution 100,000 Mishra et al. (2022)
PromptSource 100,000 Bach et al. (2022)
Math 29,837 Amini et al. (2019b)
Aqua 97,500 Ling et al. (2017)
CREAK 10,200 Onoe et al. (2021)
ESNLI 549,367 Camburu et al. (2018)
QASC 8,130 Khot et al. (2020)
SODA 1,190,000 Kim et al. (2022)
StrategyQA 2,290 Geva et al. (2021)
UnifiedSKG 779,000 Xie et al. (2022)
GSM8K 7,470 Cobbe et al. (2021)
ApexInstr 1,470,000 Huang et al. (2024b)
DeepmindMath 379,000 Saxton (2019)
DialogueStudio 1,070,000 Zhang et al. (2023)
Finance Fineweb-Fin 4,380,000 -
Book-Fin 4,500 -
Total 10,177,294
Table 3: Summary of curated texts. New datasets released with FinDaP are color-highlighted for emphasis.
For domain concept, our goal is to collect both a large volume of data and maintain high quality. Following practices from the literature on training general LLMs (Lambert et al., 2024; Gunasekar et al., 2023), we source financial texts from primarily relevant websites and books. Details on the collection process can be found in Appendix A.
Capability Domain Task IT Dataset Size Reference
Tasks Finance Relation Cls. FingptFinred 27,600 Sharma et al. (2022)
NER FingptNERCls 13,500 Yang et al. (2023)
FingptNER 511 Alvarado et al. (2015)
Headline Cls. FingptHeadline 82,200 Sinha et al. (2020)
Sentiment Cls. SentimentCls 47,600 Yang et al. (2023)
SentimentTra 76,800 Yang et al. (2023)
Summariz. TradeTheEvent 258,000 Zhou et al. (2021)
IF/Chat General IF/Chat SelfInstruct 82,000 Wang et al. (2022)
SlimOrca 518,000 Lian et al. (2023)
UltraChat 774,000 Ding et al. (2023)
ShareGPT 100,000 Link
Finance QA FinanceInstruct 178,000 Link
FingptConvfinqa 8,890 Chen et al. (2022)
FlareFinqa 6,250 Chen et al. (2021)
FlareFiqa 17,100 Yang et al. (2023)
Reasoning Math QA OrcaMath 200,000 Mitra et al. (2024)
MetaMathQA 395000 Yu et al. (2023)
MathInstruct 262,000 Xiang Yue (2023)
Code QA MagicodeInstruct 111,000 Luo et al. (2023)
Finance CFA Exam Exercise 2,950 -
Total 3,161,401
Table 4: Summary of our curated prompts. New datasets released with FinDaP are color-highlighted for emphasis. For datasets without formal references but only a URL, we provide their links.
Prompt Curation. Prompts represent the diverse ways users may interact with models and serves the essential component for IT and PA. Table 4 summarizes the prompts curation datasets. Specifically, we conduct a broad survey and source general, financial, instruction-following, and reasoning tasks from public datasets. To promote diversity, we include datasets like Flare-FinQA  (Chen et al., 2021), a large open QA dataset in finance, and UltraChat (Ding et al., 2023), a dataset shown to perform well for IT in the literature (Tunstall et al., 2024; Ivison et al., 2024). Additionally, we find that exercises or demonstrations from books that were curated in §3.1 is valuable for reasoning tasks as they usually involve challenging reasonings and come with ground truth answers and sometimes even include human-written chain-of-thought (CoT) explanations. Details on the prompts used to extract these exercises can be found in Appendix B.
3.2 Continual Pre-training (CPT)
Figure 2: Average performance on selected datasets for training Llama3-8b-instruct on our CPT-In, CPT-Gen and CPT-Mix. The selected datasets are chosen for illustration purpose based on their ability to illustrate the general trend.
In order to expose the LLM to domain-specific concepts, we first conduct continual pre-training (CPT). In CPT, we feed plain text to the LLM and perform next token prediction.
From Text to CPT Data. A key challenge in CPT is what kind of data we should use. Given the general and domain-specific texts introduced in §3.1, we can construct three versions of CPT data, CPT-In contains only the financial (in-domain) text, CPT-Gen contains only the general domain data, and CPT-Mix contains the mixture of the CPT-In and CPT-Gen.
Key Data Experiments. We conduct CPT on each of the three versions of data. As shown in Figure 2, we observe that while CPT-In and CPT-Gen outperforms in financial (Fig 2a) and general (Fig 2b) tasks, respectively, CPT-Mix achieves the best overall. This is expected as CPT-In can cause catastrophic forgetting on the general tasks, while incorporating general domain concepts in CPT-Mix acts as ‘replay’ mechanism to mitigate it (Scialom et al., 2022). We can also see that none of the CPT-trained LLMs outperform their base. This is unexpected because CPT invovles post-training on more specialized data, which should enhance the performance. By analyzing the output, we attribute this issue to the model forgetting how to follow instructions effectively after CPT. To quantify this finding, we evaluate the instruction following ability of these models using MT-Bench. The two-turn average scores for CPT-Mix, CPT-In, and CPT-Gen are 1, 1, and 1.0125, respectively, while the base model, achieves a score of 7.8875. These confirm that the conventional CPT applied to a instruction-tuned LLM can cause serious forgetting on instruction-following (IF) capability. In §3.3, we will see how jointly train IT and CPT can help mitigate such forgetting issue.
3.3 Instruction Tuning (IT)
To adapt the LLM to domain-specific and IF tasks, we conduct IT. The key different between IT and CPT is that IT masks out the instruction and takes as input supervised tasks.
From Prompt to IT Data.
Figure 3: Average performance on selected datasets for training Llama3-8b-instruct on our IT-In, IT-Gen and IT-Mix.
We introduced our prompt curation in §4. We create the responses for IT by filtering existing responses or creating new responses. For prompts with existing responses, we generally keep the original responses if they were written by a human or a strong model, such as GPT-4. We also filter out empty responses. For prompts without responses, for example, exercises extracted from books that may not have solutions provided, we generate new responses using GPT-4o. Similar to CPT data, we construct three versions of IT data, IT-In, which contains only financial (in-domain) tasks, IT-Gen, which contains only general tasks, and IT-Mix, which includes a mixture of the IT-In and IT-Gen.
Key Data Experiments. Similar to CPT, we conduct IT to each of the three versions. From Figure 3, we observe that unlike CPT, forgetting is significantly reduced. Specifically, all versions of IT are no longer worse than their base versions, indicating that the ability to follow instructions is not as severely forgotten as in CPT. This is further supported by the MT-Bench scores, where we obtained 7.2031, 6.2094, and 7.3219 for IT-Mix, IT-In and IT-Gen, repsetively, all of which are significantly better than the CPT counterparts.
Figure 4: Average performance on selected datasets for training Llama3-8b-instruct on IT-Mix with full-model finetuning (IT-Mix) and LoRA finetuning (IT-Mix (LoRA)).
We observe that IT-Mix is slightly better than other data versions, suggesting that mixing general tasks remains helpful to mitigating forgetting of general concepts and tasks, although the effect is much less pronounced compared to CPT. We also see that similar tasks improve significantly over base model while novel tasks (including financial tasks and general tasks) show little change. This indicates that, in contrast to CPT, domain has less impact in IT, but task generalization is a challenging issue.
Comparison with LoRA. Another popular approach to adapt the LLM to specific domain is Parameter-efficient Fine-tuning (PEFT), where the LLM parameters remain fixed, and only a small set of additional parameters are trained. This approach naturally mitigates forgetting issues and is more efficient in terms of trainable parameters. However, whether it can achieve performance comparable to full-model training is unclear. In Figure 4, we experiment with PEFT, specifically using LoRA (Hu et al., 2021), with a rank size of 128, [^2] Further decreasing or increasing the rank size did not show improvement in our preliminary experiments. For example, rank size of 32, 128 and 512 yield overall averages across 10 general tasks of 0.5267, 0.5331, and 0.5215, respectively, showing only minor differences. and compare its performance with full-model finetuning (IT-Mix). We observe that with and without LoRA performs similarly, confirming that LoRA is effective for task adaptation. However, the novel tasks still show little improvement, highlighting that task generalization still remains a significant challenge.
Figure 5: Average performance on selected datasets for training Llama3-8b-instruct on CPT-Mix and IT-Mix jointly (CPT-Mix + IT-Mix) and sequentially (CPT-Mix $\rightarrow$ IT-Mix).
A plausible reason for the lack of task generalization is that effective generalization may require exposure to a diverse range of tasks (Wei et al., 2022), which is often impractical in certain domains, particularly long-tail ones. However, concepts themselves may be inherently more generalizable due to the shared nature of concepts across tasks. Based on this, we propose adding CPT either before or concurrently with the IT stage and conduct training experiments accordingly.
3.4 Combining CPT and IT
A natural choice is to conduct CPT and IT sequentially (Lambert et al., 2024). On the one hand, this is flexible as it allows for different settings (e.g., data size) in each stage. On the other hand, it does not help prevent forgetting during the CPT stage, leaving the LLM dependent on IT to ‘recover’ its instruction-following capability. To make a more grounded decision, we conduct experiments on both sequential and joint training approaches. In joint training, an additional hyperparameter to consider is the mixture ratio. We find that down-sampling CPT data to match the size of IT data is the most effective strategy. Results for other strategies, such as no-sampling are provided in Appendix D.
Figure 6: Average performance on selected datasets for PEFT or full model fine-tuning for CPT and IT.
Figure 5 illustrates the comparison between joint and sequential training. In both cases, different from IT-only results shown in Figure 3, we see improved performance on similar and novel tasks. This supports our hypothesis that CPT can help improve the generalization of IT, as the concepts are likely able to be shared across different tasks. It is further interesting to see that even the general tasks are improved, indicating that there could be positive transfer between CPT and IT. Comparing the two, we observe that joint training outperforms sequential training across financial and general tasks, as well as similar and novel tasks, highlighting the importance of preventing forgetting of CPT and knowledge transfer between CPT and IT.
Comparison with LoRA. In Section 3.3, we showed that LoRA can effectively adapt tasks but still suffers from task generalization. While we already showed that CPT can help in full-model training setting, we now explore whether CPT can help in the PEFT setting as well. Figure 6 presents the results of applying LoRA for IT and LoRA for both CPT and IT. Surprisingly, we find that full fine-tuning significantly outperforms the LoRA counterparts across similar and novel tasks. This finding contrasts with our previous observations in Figure 4, where performance with and without LoRA was comparable. Our results reveal that knowledge transfer from CPT to IT, which is crucial for task generalization, requires full-model training.
Takeaways from Continual Pre-training and Instruction Tuning CPT and IT have different effects and challenges. CPT is effective at introducing domain concepts. However, addressing the forgetting of general concepts and IF capability is crucial. In contrast, IT is effective at introducing domain-specific tasks, with negligible forgetting of IF capabilities and only slight forgetting in general tasks. To improve generalizability, mixing general-domain text and general tasks, along with joint training for IT and CPT is found to be more effective. Furthermore, PEFT is beneficial for task adaptation, but transferring knowledge from CPT to IT to improve task generalization requires full-model training.
3.5 Improving Reasoning with Preference Learning
In §3.2 & §3.3, we observed that with CPT and IT, the model improves in capabilities such as concepts, tasks and IF/Chat. However, such improvements were missing on the reasoning tasks. Preference Alignment (PA), where the model is trained to assign higher probability mass to better generations, has been shown to be effective in enhancing reasoning capabilities of LLMs (Pang et al., 2024; Lambert et al., 2024; Jiao et al., 2024; Wang et al., 2024). Specifically, we employ Direct Preference Optimization (DPO)  (Rafailov et al., 2023), which directly learns from positive (chosen) and negative (rejected) preference data, yielding an easier way for PA compared to reinforcement learning. We synthetically generate such data from the on-policy model, i.e., the jointly trained CPT+IT checkpoint, as it has shown the strongest performance in previous experiments.
Figure 7: Average performance on selected datasets for PA training from the ‘CPT+IT’ checkpoint on PA-Mix and PA-In.
Negligible Forgetting in PA. As with CPT and IT, we begin by performing an ablation study on different data versions to evaluate their effectiveness. Since the degree of forgetting diminishes from CPT to IT (as observed in §3.3), we expect it to be even less pronounced in PA. To quickly evaluate this hypothesis, we take a naive approach and create PA-Mix and PA-In by using either the provided or GPT4o generated responses (as done for IT in §3.3) as the ‘chosen’ samples and the output of ‘CPT+IT’ checkpoint as the ‘rejected’ ones, based on the prompts of IT-Mix and IT-In, respectively.
Figure 7 shows the results after PA training for PA-In and PA-Mix from the ‘CPT+IT’ checkpoint. [^3] PA trained from Llama3-8b-instruction has shown worse results compared to training from the ‘CPT+IT’ checkpoint in our preliminary experiments, as PA requires a strong initialization checkpoint. For instance, PA-Mix from Llama3-8b-instruction achieves only 29.99 on EDTSUM, whereas the CPT+IT’ counterpart achieves 54.21. As a results, we only investigate training PA from ‘CPT+IT’ checkpoint. We observe that PA-In performs comparably to PA-Mix, indicating that it may not be essential to include general tasks to prevent forgetting of concepts or tasks, unlike the cases of CPT and IT. This suggests that PA training can focus on in-domain tasks, without requiring a broader set of general tasks or raising concerns about forgetting. Given this, we use CFA exams (Table 4 in §3.1) as a representative source for in-domain reasoning because they cover diverse financial scenarios, emphasize complex reasoning, and, most importantly, are derived from real-world exams. These characteristics make them a strong proxy for a broader range of financial tasks, ensuring that the model generalizes effectively within the financial domain while simplifying the training process.
Another crucial observation is that there is not much difference even for unseen similar tasks (FiQA SA, FOMC and NER) and reasoning tasks (CFA-Easy and CFA-Challenge). This highlights the limitations of the current naive PA approach and suggests room for further improvement. In the following, we propose a novel PA approach that constructs preference data guided by both outcome and process reward signals.
Proposed Method and Comparison with Prior Work. Substantial efforts have been made in the past to improve reasoning with LLMs. These lead to two main families of approaches: inference-time (Snell et al., 2024; Wu et al., 2024) and training-time (Jiao et al., 2024; Qu et al., 2024) methods. Inference-time methods such as search-based (Snell et al., 2024) and self-correction (Qu et al., 2024) achieve good performance by scaling up inference-time computation. Despite their effectiveness, these methods often suffer from high latency due to the frequent assessment of intermediate reasoning states and the extensive exploration space. For example, a Monte Carlo Tree Search (MCTS) method (Jones, 2021) may improve task performance, but potentially requires significantly more computation than simply sampling solutions multiple times. This makes them unsuitable for the finance domain, where applications often require rapid responses, such as for real-time market analysis or portfolio management. In contrast, training-time methods first collect trajectories using (variations of) inference-time methods and then train the LLM with the collected trajectories. This helps the model reason more accurately and faster during inference. Therefore, in this work, we adopt the training-time approach and limit ourselves to simple CoT based inference methods.
Figure 8: An overview of the proposed Trajectory-GenORM and Trajectory-GenPRM. In Trajectory-GenORM, we collect trajectories from GenORM by evaluating the entire solution (thus question marks in the intermediate steps). In Trajectory-GenPRM, we collect trajectories from GenPRM, by identifing and correcting the first erroneous step.
To collect reasoning trajectories, there are two main approaches. The first is search-based (Setlur et al., 2024; Snell et al., 2024), where a trained Reward Model (RM) or verifier is used to guide a search method (e.g., Best-of-N, Beam Search) to identify the best reasoning path. The second is revision-based (Bai et al., 2022; Du et al., 2023; Madaan et al., 2023; Saunders et al., 2022), which attempts to improve the generation distribution through multi-round interactions, often by leveraging feedback from itself or another strong LLM to refine the input prompt.
In practice, revision-based methods have shown mixed results and have not yet been well established as reliable for achieving improvements (Huang et al., 2024a). In contrast, search-based methods have been shown to be more effective but they heavily rely on the performance of the RM. Most RMs typically assign numerical scores to candidate solutions, which do not fully exploit the text-generation capabilities that LLMs are inherently designed for. Consequently, there has been growing interest in Generative Reward Models (GenRMs), where the RM is a generative model capable of producing text-based feedback. However, most existing GenRM still either provide only outcome-level (i.e., on the final answer) signals (Mahan et al., 2024) or require extensive training (Zhang et al., 2024), which results in sparse supervision and under-utilizes the strengths of pre-trained LLMs.
To overcome these limitations, we propose leveraging the advanced text generation capabilities of LLMs as a GenRM to provide rewards on both the final answer (outcome reward) and the reasoning steps (process reward). Figure 8 illustrates the proposed method. Specifically, we construct preference data from Generative Outcome Reward Model (GenORM) and Generative Process Reward Model (GenPRM) trajectories, both of which are implemented using GPT-4o:
$\bullet$ Trajectory from GenORM (Trajectory-GenORM). Given a prompt and a candidate model generated solution, we prompt the GenORM to give a holistic judgment for the entire solution using a single ‘Yes’ or ‘No’ token. We then use the correct solutions as chosen samples and the incorrect solutions as rejected samples.
$\bullet$ Trajectory from GenPRM (Trajectory-GenPRM). Since reasoning is often complex and process rewards have been shown to be more effective (Lightman et al., 2024), we further leverage GenPRM to provide process rewards. Instead of requesting rewards at each step, which has been shown to be unnecessary by Lightman et al. (2024) and Luo et al. (2024), we prompt the GenPRM to identify the first erroneous step. While identifying the first error is helpful for training a Reward Model (RM), it is insufficient for generating preference data. To address this, we further extend the approach by prompting the GenPRM to correct the identified erroneous step.
Figure 9: Average performance on selected datasets for training from ‘CPT+IT’ on PA using Trajectory-GenORM, noted as PA (GenORM) and Trajectory-GenORM + Trajectory-GenPRM, noted as PA (GenORM+GenPRM).
Using this correction, we construct a preference data sample. The input prompt is formed by concatenating the original question, the candidate reasoning steps up to the first error, and a follow-up question framed as “What is the next step?”. The chosen response of this preference sample is the newly-obtained corrected step, while the original first erroneous step is deemed as rejected response. This approach produces trajectories that focus on predicting the correct next step given a reasoning prefix, rather than requiring a prediction of the entire reasoning trajectory, as in Trajectory-GenORM. Additional details about the prompts used for our data collection are provided in Appendix C.
Results of the Proposed Method. Figure 9 shows the results. We observe significant improvement in both CFA-Easy and CFA-Challenge, indicating that the supervision provided by Trajectory-GenORM and Trajectory-GenPRM is highly beneficial for reasoning tasks. Moreover, combining Trajectory-GenORM and Trajectory-GenPRM yields additional improvements over Trajectory-GenORM alone, demonstrating that outcome and process rewards complement each other effectively.
Takeaways from Preference Alignment PA is effective for learning to reason. Similar to IT, the forgetting of general concepts and IF capability is negligible, as is the forgetting of general tasks. However, synthesizing preference data for effective adaptation is challenging. Constructing preference data by collecting trajectories from both outcome and process signals of a GenRM is found to be an effective method.
4 Final Recipe and Evaluation
In this section, we present our final recipe based on the desired capabilities in Section 2.1, and data and model ablations performed in Section 3. A summary of the final recipe is given in Table 5. For evaluation, we compare with a wide range of baselines models, including Llama-Fin against a wide range of baseline models, including its base model, Llama3-8B-instruct, and the 8B peer, Llama3.1-8B-instruct. We also include comparisons with models of other sizes, such as Phi-3.5-mini-instruct (Abdin et al., 2024) (3.8B), and Mistral-Nemo-instruct (Jiang et al., 2023) (12B), as well as the closed model GPT-4o (OpenAI, 2023). Additionally, we evaluate against the latest open finance-specific LLM, Palmyra-Fin-32k (Writer, 2024), which is also based on the Llama3-8B-instruct model. [^4] We note that there are additional financial LLMs available, such as FinMa (Xie et al., 2023) based on Llama2, Finance-LLM (Cheng et al., 2024) based on Llama3-base and FinLLaVA (Xie et al., 2024b) focuses on multi-modality and not publicly available. However, they are either significantly smaller in scale, based on less advanced LLMs compared to our model or not publicly available. In our preliminary experiments, these models performed considerably worse than both our model and the baselines. Therefore, we have only included the SoTA financial LLM in our comparisons.
Final Recipe for Llama-Fin
Continual Pre-training (CPT) and Instruction Tuning (IT)
Data 50% CPT, 50% IT
Curriculum Group 1 CPT: 50% Domain-specific Text (Web and book), 50% General text (verfiable text)
IT: 20% Domain-specific tasks, 80% General tasks
Group 2 CPT: Group 1 data + domain-specific books
IT: Group1 + Exercises extracted from books
Steps
Group 1: 3.84B tokens; Group 2: 1.66B tokens
(8,000 context length, 16 A100)
Model Intialization Llama3-8b-instruct
Attention CPT: full attention with cross-docuemnt attention masking
IT: full attention with instruction mask-out and cross-docuemnt attention masking
Optim. AdamW (weight decay = 0.1, $\beta_{1}$=0.9, $\beta_{2}$=0.95)
LR Group 1: 5e-6 with 10% warmup; Group 2: 5e-6 with 50% warmup
Batch size 128K tokens
Stop Cri. Loss of development set stops decreasing ($\approx$ 1 epoch)
Preference Alignment (PA)
Data Trajectory-GenORM and Trajectory-GenPRM
Steps 24.58 M tokens
Model Initialization CPT+IT
Loss DPO with an additional negative log-likelihood term
Optim. LR 5e-7 with 10% warmup
Batch size 32K tokens
Stop Cri. Loss of development set stops decreasing
Table 5: Final recipe of Llama-Fin. The joint training of CPT and IT is structured into two groups, with each group undergoing joint training sequentially. The second group utilizes higher-quality data (sourced from books), following the typical curriculum training practice (Gao et al., 2024). For PA, we employ a modified DPO loss with an additional negative log-likelihood term, similar to Pang et al. (2024), as it has shown to be more effective than relying solely on the original DPO loss.
Capability Domain Task Benchmark
Llama-Fin
8B
CPT+IT
8B
Llama3
Instruct
8B
Llama3.1
Instruct
8B
Palmyra
Fin
70B
Phi
3.5-mini
Instruct
3.8B
Mistral
Nemo
instruct
12B
GPT4o
Unseen - Similar
Tasks Finance Sentiment Analysis FPB (Acc) 91.13+ 92.99 73.09 71.55 67.11 78.04 78.25 82.16
Sentiment Analysis FiQA SA (Acc) 95.32+ 94.47 77.87 70.64 71.91 69.36 55.74 68.51
Monetary Policy Stance FOMC (Acc) 64.31+ 63.10 56.65 54.64 63.10 58.47 57.86 67.94
Named Entity Recognition NER (Rouge1) 76.69+ 74.33 45.03 51.22 54.29 39.37 49.84 43.02
Abstractive Summarization EDTSUM (Rouge1) 53.78+ 54.21 11.50 12.53 21.77 19.97 12.32 18.15
Unseen - Novel
Concept General Knowledge Recall MMLU (CoT, Acc) 47.42 47.22 48.14 47.42 54.93 45.07 49.64 63.88
AI2-ARC (CoT, Acc) 89.43+ 88.95 89.29 89.80 89.01 87.25 88.19 97.85
Nq-open (CoT, Acc) 19.20+ 16.20 18.47 22.52 19.25 6.20 17.01 27.92
Finance Knowledge Recall MMLU-Finance (Acc) 64.20 63.93 65.71 66.74 75.15 68.17 61.88 86.52
Task Finance Extractive Summarization Flare-ECTSUM (Rouge1) 34.10 34.41 35.92 35.77 33.24 35.52 37.86 35.90
ESG Issue Classification MLESG (Acc) 40.67+ 42.00 36.33 36.00 39.67 38.33 32.67 45.67
Rumor Detection MA (Acc) 84.00+ 84.60 82.60 84.20 62.60 75.40 85.20 73.80
Stock Movement Prediction SM-Bigdata (CoT, Acc) 54.14 52.04 55.3 46.06 48.70 53.26 53.53 49.18
SM-ACL (CoT, Acc) 51.99+ 49.89 50.51 45.30 51.21 49.84 50.75 50.97
SM-CIKM (CoT, Acc) 54.94 44.88 55.56 48.03 52.92 50.03 53.28 49.78
Fraud Detection CRA-CCF (CoT, Mcc) 0.83+ 0.61 -0.32 2.73 3.12 1.20 3.94 6.16
CRA-CCFraud (CoT, Acc) 34.03+ 32.32 14.78 17.3 33.03 45.33 32.94 49.57
Credit Scoring Flare-German (CoT, Acc) 64.00+ 60.50 33.50 15.00 12.00 49.50 32.50 17.00
Flare-Astralian (CoT, Acc) 44.60 51.80 66.91 11.51 12.95 46.76 56.12 51.80
CRA-LendingClub (CoT, Acc) 68.49+ 65.96 52.69 25.38 23.40 48.87 21.03 65.03
Distress Identification CRA-Polish (CoT, Mcc) 15.30+ 0.65 12.37 15.07 13.78 69.14 11.18 17.38
CRA-Taiwan (CoT, Acc) 40.81+ 96.41 12.01 35.97 52.58 69.96 57.88 8.57
Claim Analysis CRA-ProroSeguro (CoT, Acc) 35.14 86.57 96.98 44.33 56.20 25.86 32.58 96.60
CRA-TravelInsurance (CoT,Acc) 41.52+ 98.50 6.39 80.31 17.28 94.48 73.64 54.03
Tabular QA *Flare-TATQA (CoT, Acc) 66.61+ 66.43 67.80 63.70 64.21 57.70 66.40 74.90
Open QA *Finance Bench (CoT, Acc) 54.00+ 52.00 52.70 38.00 56.67 40.70 55.30 51.30
IF/Chat General Precise IF MT-bench (1,2 turn avg) 7.36 7.29 7.88 7.92 5.80 8.38 7.84 9.10
Reasoning Math Math Reasoning MathQA (CoT, Acc) 55.08+ 54.30 51.16 49.35 41.51 39.40 52.46 70.82
General Social Reasoning Social-IQA (CoT, Acc) 75.23+ 73.64 68.83 70.73 77.28 72.82 62.95 78.92
Common Sense Reasoning Open-book-qa (CoT, Acc) 82.60+ 79.20 77.00 82.20 87.00 80.20 76.40 94.60
Hellaswag (CoT, Acc) 81.90+ 78.92 73.34 69.10 69.69 67.89 61.74 81.76
Winogrande (CoT, Acc) 70.32+ 67.48 62.51 66.69 74.27 72.22 65.82 85.71
PIQA (CoT, Acc) 85.85+ 84.39 79.82 81.45 86.72 82.05 77.91 94.34
Finance Exam CFA-Easy (CoT, Acc) 66.28+ 62.31 60.56 60.47 36.05 61.24 65.89 83.14
CFA-Challnge (CoT, Acc) 55.56+ 35.56 34.44 35.56 25.56 48.89 43.33 74.44
Table 6: Overview of the results on unseen evaluation set. ‘*’ indicates that ‘GPT4o’ is used as the judge. Llama-Fin and its variant (‘CPT+IT’) are highlighted in blue. The best performing model for 8b on each benchmark is bolded. The overall best performance across all models is underlined. + indicates that Llama-Fin outperforms the base Llama3-8B-instruct.
Table 6 shows the evaluation of Llama-Fin across all considered benchmarks. Scores are reported using the metrics specified in parentheses555‘Acc’ and ’Rouge1’ refer to accuracy and ROUGE-1 score respectively. ‘Mcc’ refers to Matthews correlation coefficient, usually used in highly imbalanced data (Xie et al., 2024b)., ensuring consistency with the corresponding literature. Higher scores always indicate better performance. It is important to note that while all evaluations are unseen for Llama-Fin, we cannot verify whether some benchmarks were used during the training of the baseline models.
Unseen - Similar. Llama-Fin trained based on Llama-3-8b-instruct outperforms all other baselines in its size category. It also surpasses significantly larger models, such as the finance-specific Palmyra-Fin-32K (70B). Notably, Llama-Fin also exceeds the performance of the closed model GPT-4o. These results demonstrate the effectiveness of our data and model recipe for domain-adaptive post-training.
Unseen - Novel. To evaluate the generalization of Llama-Fin, we evaluate its performance on unseen novel tasks that correspond to the identified capabilities. Below, we summarize the key takeaways from this comparison:
$\bullet$ Llama-Fin Preserves General Concepts. We observe that Llama-Fin performs better or remains competitive with its base model in general knowledge recall tasks, indicating that it effectively preserves general concepts. It performs slightly worse than the base model in finance knowledge recall (MMLU-Finance), despite our earlier finding that the CPT benefits IT (§3.3). We hypothesize that CPT helps learn concepts that are helpful but differ from those emphasized in MMLU-Finance.
$\bullet$ Llama-Fin is Effective in The Majority of Tasks. Llama-Fin outperforms the base model in 12 out of 17 tasks, demonstrating that our approach can lead to models that generalize well to novel, unseen tasks requiring the same capabilities. Interestingly, we observe that for certain tasks (e.g., CRA-TravelInsurance, CRA-Taiwan and CRA-ProroSeguro), preference alignment negatively impacts performance, resulting in worse outcomes compared to ‘CPT+IT’. Even GPT-4o performs poorly in these tasks. This suggests that for some tasks, leveraging reasoning capabilities might not be beneficial, as these tasks could be inherently ‘easy’ and solvable without the need for advanced reasoning. Such observations aligns with findings in the literature (Sprague et al., 2024; Liu et al., 2024).
$\bullet$ Llama-Fin Preserves IF/Chat Capabilities. Llama-Fin achieves a competitive MT-Bench score compared to the base model, indicating that it effectively maintains the IF capability.
$\bullet$ Llama-Fin Excels in Reasoning Tasks. For reasoning capability, Llama-Fin significantly outperforms the base models across all considered benchmarks in a large amount (up to 20% in CFA-Challenge), indicating substantial improvements in reasoning capability. Moreover, Llama-Fin is significantly better than ‘CPT+IT’, further confirming that our proposed Trajectory-GenPRM and Trajectory-GenORM are particularly effective in improving reasoning performance beyond the already strong checkpoint of ‘CPT+IT’.
5 Conclusion
We introduce FinDaP, an open SoTA finance-specific post-training framework, featuring evaluation framework, leaderboard, data, and model recipes. Under FinDaP, we develop Llama-Fin, a SoTA finance LLM. In this development, we conduct a systematic study on effectively adapting a target domain through post-training. This includes identifying key capabilities, designing evaluation setups, and detailing training and data recipe across the stages of CPT, IT, and PA. Furthermore, we propose a novel approach for constructing preference data during PA. For each stage, we reveal the distinct challenges, objectives, and effective strategies. Looking ahead, we aim to explore additional domain-specific capabilities using FinDaP.
6 Limitations
While the recipe for FinDaP and Llama-Fin are effective, the performance on novel unseen tasks still requires further improvement. For example, selectively employing reasoning capabilities only for questions that require such advanced reasoning might give better results. Additionally, the data recipe is currently based on full-scale empirical experiments, which can be time-intensive. Developing low-cost experiments to reliably indicate the effectiveness of data in post-training could streamline this process and accelerate the development iteration. It is also worth noting that the same recipe may not generalize well to other model families. Different architectures or pretraining strategies might require tailored recipe to achieve optimal results, emphasizing the need for adaptability in recipe design in future research.
Appendix A Texts Curation Details
we source financial text from two primary resources. The first source is web text, where we filter non-financial content from the FineWeb using URLs like ‘sec.gov’ and ‘investopedia.com’. The second source is books. We manually select 10 finance-related topics (e.g., ‘economics’ and ‘management’), download books on these topics, and convert them to text using OCR (Malmgren, 2014). Since OCR can make mistakes, we further employ a strong LLM to filter out content lacking educational value or unrelated to finance. Details on the financial URLs, finance-related topics, and the prompts used for filtering is shown below:
Selected Financial URLs. We curated a selection of 70 financial websites to comprehensively cover diverse aspects of finance-related content on the web. These include trusted sources from financial institutions, regulatory agencies, educational platforms, and industry-specific news outlets. This diverse collection ensures representation across sub-domains such as investment, banking, personal finance, regulatory compliance, and financial planning, offering a well-rounded foundation that can cover most of the finance content in the web.
Selected Topics. We select 12 topics that are cover most of books in finance. 5 of them are from business areas, including business, Accounting, Accounting, Management, Marketing, Trading; 1 is from Mathematics, i.e., Mathematical Economics; and 4 are from Economy area, including economy, econometrics, investing, and markets. We crawled the web to downloaded the books from the corresponding topics. For CFA, we use the material provided by CFA prep providers.
Prompt for Filtering the Text. We explored various prompt formats to automatically extract an financial score using an LLM and found that the additive scale by Yuan et al. (2024) worked best. Figure A.1 shows the prompt we used to filter the ‘low-quality’ text. Specifically, this prompt allows the LLM to reason about each additional point awarded, unlike the single-rating Likert scale which fits samples into predefined boxes. Then, to avoid the LLM favoring highly technical content like academia papers, we focused on financial student level knowledge. By setting a threshold of 4 (on a scale of 0 to 5) during the filtering process, we were able to also retain some high-quality financial content.
Prompt for Filtering the Text Below is an extract from a text book. Evaluate whether the book has a high financial value and could be useful in an financial setting for teaching financial students using the additive scoring system described below. Points are accumulated strictly based on the satisfaction of each criterion: - Add 1 point if the extract provides educational value for financial students whose goal is to learn financial concepts or take finacial exams. It is acceptable if quizzes are not included; however, if quizzes are present, detailed solutions and explanations must also be provided Add 1 point if the extract provides educational value for financial students whose goal is to learn financial concepts or take financial exams. It is acceptable if quizzes are not included; however, if quizzes are present, detailed solutions and explanations must also be provided. Add another point if the extract addresses certain elements pertinent to finance and aligns closely with financial standards. It might offer a superficial overview of potentially useful topics or present information in a disorganized manner and incoherent writing style. Award a third point if the extract is appropriate for financial use and introduces key concepts relevant to financial curricula. It is coherent and comprehensive. Grant a fourth point if the extract is highly relevant and beneficial for financial learning purposes for a level not higher than financial students, exhibiting a clear and consistent writing style. It offers substantial financial content, including exercises and solutions, with minimal irrelevant information, and the concepts aren’t too advanced for financial students. The content is coherent, focused, and valuable for structured learning. Bestow a fifth point if the extract is outstanding in its financial value, perfectly suited for teaching either at financial students. It follows detailed reasoning, the writing style is easy to follow and offers profound and thorough insights into the subject matter, devoid of any non-financial or complex content. The extract: <EXAMPLE>. After examining the extract, You will output a json object containing the following 2 fields: { "Justification": string // Briefly justify your total score, up to 100 words. "Score": integer // Conclude with the score }
Figure A.1: Prompt for filtering the text
Appendix B Prompts Curation Details
Figure B.1 shows the prompt we used to extract exercises from books. We carefully design the prompt to extract both the question part of an exercise, which potentially include questions, scenario and exhibits, and the answer part of the exercise, which may include answer choices and solution. In books, questions and their corresponding answers can be located far apart (e.g., the questions may appear at the beginning while the solutions are provided at the end), meaning they may not be captured within the same chunk. As a result, some questions may not have corresponding extracted answers. For such cases, GPT-4o’s generated answers are used when converting the prompt into instruction-following or preference-alignment data, as detailed in Sections 3.3 and 3.5.
Prompt for Extracting Exercise from Book You are an educational assistant aims to extract all questions from the provided material. Look for specific indicators such as "example," "quiz," "questions," or similar terms to identify where the questions are located. If the material includes scenarios or exhibits, must include all details related to them. Do not create or derive any questions or come up with content on your own—strictly extract what is present in the material. Make sure no question is missed. If one scenario or exhibits corresponds to multiple questions, duplicate the scenarios and exhitbits so that the number of questions match the number of scenarios and exhibits. The material: <MATERIAL>. After performing these tasks, You will output a json object containing the following fields: { "Justification": "string", // A brief justification for your extractions, up to 100 words. "Questions": "string", // A list of questions extracted from the material. Only extract the exact questions presented in the text. "Scenario": "string", // A list of scenarios corresponding to the above questions. If the material does not provide the scenario place "N/A." Do not do any derivation or reference, must output the exact same, detailed and complete scenarios. The scenario may contain multiple paragraphs or even splited by the exhibits, combine them into one string. The scenario can be long, you may modify it to make it shorter, but must not change its meaning. "Exhibit": "string", // A list of exhibits or tables corresponding to the above questions. If the material does not provide the exhibit, place "N/A." Do not do any summary, or derivation or cutting, must output the exact same, detailed and complete exibits. There may be multiple exhibits involved in a scenario, combine them into one string. The exhibit can be long, you may modify it to make it shorter. Must keep the table format "Answer Choices": "string", // A list of answer choices corresponding to the above questions. If the material does not provide answer choices, place "N/A." "Answer": "string" // A list of answers corresponding to the above questions. Answers should only be included if provided in the material. If no answer is given, place "N/A." If explanations or reasoning steps or equations are included, must capture all of them. Must not answer it yourself if there is no answer provided in the material. Make sure the final number of questions equals to number of scenario equals to number of exhibits equals to number of answers }
Figure B.1: Prompt for extracting exercises from books
Appendix C GenRM Details
In §3.5, we simplified the prompt for GenRM for the purpose of illustration in Figure 8. In this section, we give full detailed of the prompt for GenORM and GenPRM in Figure C.1 and Figure C.2, respectively.
Prompt for GenORM You given a question, a reference answer and a proposed answer, you task is to determine the correctness of the proposed answer. First, extract the final answer (for example, A, B or C) from the reference answer. Second, extract the final answer from the proposed answer (for example, A, B or C). Finally, compare the two final answer to determine the correctness. Do not do any extra reasoning, must determine the correctness soley based on the given reference and proposed answer. Question: <QUESTION>
Reference Answer: <REFERENCE>
Proposed Answer: <PROPOSAL>
After performing these tasks, You will output a json object containing the following fields: { "Justification": "string", // A brief justification for your output, up to 100 words. "Correctness": "string", // If the proposed answer has the same final final answer as the reference answer (for example, both choose A or have the same answer), output correct’. Put wrong to all other cases. For example, if the proposed answer has a different final answer comparing to the reference answer, put wrong’. If the proposed answer does not explicitly give a final answer to the question, put wrong’. If the proposed answer gives more than one final answer to the question, put wrong’. }
Figure C.1: Prompt for GenORM
Prompt for GenPRM Given a question, a reference answer and an incorrect answer, you task is to identify the first incorrect step from the incorrect answer. The "first incorrect step" means all reasoning up to that point is accurate, but the error begins at this specific step. Question: <QUESTION>
Reference Answer: <REFERENCE>
Incorrect Answer: <INCORRECT>
After performing these tasks, You will output a json object containing the following fields: { "Justification": "string", // A brief justification for your output, up to 100 words. You need to explain (1) why the identified first incorrect step is incorrect; (2) why the reasoning up to this specific step is correct and (3) how the corrected step resolves the issue, aligning with the reference answer, maintaining the logical flow and progressing to the final answer. "First incorrect step": "string", // The explanation in the incorrect answer consists of multiple reasoning steps. Please identify the first incorrect reasoning step. It should be a piece of text directly and exactly quoted from the incorrect answer. It should be an intermediate step rather than the final answer "Reasoning up to incorrect": "string", // From the incorrect answer, give the correct reasoning steps up to the first incorrect step. This should be directly and exactly quoted from the incorrect answer. "Step correction": "string", // Replace the identified incorrect step with a single, clear, and correct step. This step should directly address and correct the error, explicitly providing the correct reasoning without requring for more information or challenging the question. It should effectively answer the question, "What is the next reasoning step?" given on the question and the identied "Reasoning up tp incorrect". It should help progress to the final answer. }
Figure C.2: Prompt for GenPRM
Appendix D Additional Ablations
In §3.3, we employ down-sampling CPT data to match the size of IT data when conducting CPT and IT jointly. In this section, we also evaluate the performance under a ‘no-sampling’ setting. Figure D.1 shows the results. We observe that in both joint and sequential training, down-sampling yields better results on financial tasks. This is understandable because down-sampling assigns more weight to IT, which is beneficial for the financial tasks. Interestingly, we observe the opposite trend for general tasks: no-sampling performs better. We hypothesize that this is because having more CPT data helps preserve general concepts more effectively, although it may diminish instruction-following abilities.
Figure D.1: Average and selected datasets performance from down-sampling or no-sampling on CPT.