### v2

# Domain-adaptive Post-training of Large Language Models: An Empirical Study with Llama for Finance [Yifei: A more attractive/informative title?]

Abstract
Domain-adaptive post-training of large language models (LLMs) has emerged as a promising approach for specialized domains such as medicine and finance. However, significant challenges remain in identifying optimal adaptation criteria and training strategies across varying data and model configurations. To address these challenges, we introduce FinDaP, a systematic and fine-grained investigation into domain-adaptive post-training of LLMs for the finance domain. Our approach consists of four key components: FinCap, which defines the core capabilities required for the target domain; FinRec, an effective training recipe that jointly optimizes continual pre-training and instruction-following, along with a novel preference data distillation method leveraging process signals from a generative reward model; FinTrain, a curated set of training datasets supporting FinRec; and FinEval, a comprehensive evaluation suite aligned with FinCap. The resulting model, Llama-Fin, achieves state-of-the-art performance across a wide range of financial tasks. Our analysis also highlights how each post-training stage contributes to distinct capabilities, uncovering specific challenges and effective solutions, providing valuable insights for domain adaptation of LLMs.
Demystifying Domain-adaptive Post-training for Financial LLMs
Zixuan Ke, Yifei Ming, Xuan-Phi Nguyen, Caiming Xiong and Shafiq Joty Salesforce AI Research {zixuan.ke,yifei.ming,xnguyen,cxiong,sjoty}@salesforce.com [Uncaptioned image] Project Page: https://github.com/SalesforceAIResearch/FinDAP [Uncaptioned image] Datasets: https://huggingface.co/datasets/Salesforce/FinEval
1 Introduction
While LLMs have demonstrated strong generalization across a variety of tasks, they often struggle to perform well in specialized domains such as finance and law. Consequently, domain-adaptive post-training of LLMs has garnered significant attention recently (Colombo et al., 2024a; Xie et al., 2024b). In the earlier days of language models, continual pre-training (CPT) was the dominant strategy. This involved further training a pre-trained model on domain-specific plain text and then fine-tuning it for individual tasks (Gururangan et al., 2020; Ke et al., 2023). With LLMs, the post-training focus has shifted to zero- and few-shot task generalization through methods such as instruction-tunning (IT) (aka. supervised fine-tuning or SFT) and preference alignment (PA). While prompt engineering of powerful general LLMs with zero- or few-shot examples has emerged as a convenient approach to adapting them to new tasks, to get the most optimal performance on a target domain, recent methods explore fine-tuning model wights to make them domain experts Chen et al. (2023b); Li et al. (2023); Colombo et al. (2024b).
Building on this trend, this work focuses on adapting LLMs to specific domains through parameter training. It complements semi-parametric methods that leverage external knowledge, such as retrieval-augmented generation (Lewis et al., 2020; Ke et al., 2024). Our focus is also different from general post-training, as the goal is not to develop another general-purpose LLM but to create specialized, expert-level LLMs tailored to a specific domain. By focusing on a specific domain, we develop models that are not only more compact in size but also deliver significantly more accurate and contextually relevant responses compared to general-purpose LLMs. Their smaller size enhances efficiency, optimizing both computational resource usage and training time.
Despite the potential of domain-specific LLMs, there is still no systematic study on what makes a good domain-specific LLM. In this work, we consider finance as the domain of interest and aim to address the following research questions:
Given a strong general-purpose LLM (e.g., Llama3-8b-inst), how to effectively adapt it to a target domain (e.g., finance) by post-training? What criteria are desirable for successful adaptation? What are effective training recipes with respect to data and model?
Figure 1: An overview of our finance-specific post-training framework, FinDaP. It comprises four key components: (1) FinCap, the core expected capabilities, including concepts, reasoning, instruction-following and tasks; (2) FinRec, encompassing both data and model strategies to guide domain-adaptive post-training; (3) FinTrain, which curates training texts and prompts based on the data recipe; and (4) FinEval, a comprehensive evaluation framework designed to assess performance on unseen tasks, categorized into similar and novel, general and domain-specific, and reasoning tasks, using both direct-answer and chain-of-thought (CoT) evaluation methods.
Prior studies Bhatia et al. (2024); Xie et al. (2024a) typically adopt a simplified and informal framework (see §2) in that they evaluate only on a set of domain-specific end tasks such as sentiment analysis and NER, and they simply follow standard post-training stages (CPT, IT and/or PA) without considering their impact or optimizing their recipe for domain-adaptive post-training. This simplified approach can misalign with our broader expectations for a domain-expert LLM. A domain-expert LLM should not only excel at such end tasks but also achieve broader capabilities, such as follow task instructions effectively and reason in a way that aligns with domain-specific knowledge, while retaining general capabilities.
We argue that domain-adaptive post-training poses unique challenges compared to pre-training or general post-training. There are multiple factors to be considered: (1) For a particular target domain, it is essential to establish the desirable capabilities that a domain-expert LLM should possess, as these capabilities serve as a guiding framework for the entire adaptation process; (2) The training recipe should be tailored specifically to adapt an already trained LLM (e.g., Llama3-inst) through post-training. This differs from training a model from scratch or from a base pre-trained checkpoint, as it requires careful consideration of catastrophic forgetting and knowledge transfer from the original LLM, which already possesses strong general knowledge and instruction following capabilities. Each of the standard CPT, IT, and PA stages have different impacts and trade-offs with respect to knowledge forgetting and transfer, as do the in-domain, general-domain datasets, and the mixture of them. Moreover, it should also be designed to support the desired capabilities. For example, improving reasoning capability might require more dense supervision than the final answer level correctness score. (3) The desired quantity and quality of training datasets should be carefully balanced: high-quality general-domain data is required to mitigate forgetting, while diverse data and supervision signals are necessary to learn domain knowledge. (4) Finally, the evaluation methods should align with the desired capabilities. Different evaluation techniques may be required for certain capabilities; for example, chain-of-thought (CoT) Wei et al. (2023) reasoning is often necessary to effectively evaluate reasoning tasks.
In this work, we introduce FinDaP (Figure 1), a novel finance-specific framework designed to incorporated all these factors in domain-specific post-training. To our knowledge, none of the prior studies consider all of them to provide a principled guidance on domain-adaptive post-training. FinDaP integrates four key components: (1) FinCap, a set of core capabilities required for the domain expert LLM, including domain concepts, tasks, instruction following and reasoning; (2) FinRec, a training recipe that jointly performs CPT and IT, and subsequently conducts PA, balancing trade-offs across these stages to mitigate forgetting and improve task generalization. It also proposes to use mixture of in-domain and general domain data in the data recipe, alongside a novel preference alignment method for improving reasoning capability that constructs data using the preference signal in reasoning steps, Stepwise Corrective Preference (SCP) and final answer, Final Answer Preference (FAP); (3) FinTrain, a curated set of training datasets implementing FinRec, which carefully balances quality and diversity; and (4) FinEval, a comprehensive evaluation framework covering a wide range of tasks, including reasoning tasks assessed through CoT.
We apply FinDaP on the instruction-tuned Llama3-8b-instruct (LLaMA, 2024). Our best performing recipe yields Llama-Fin that outperforms all considered baselines, including large open models at the 70B scale and proprietary models like GPT-4o, on tasks that are similar (yet unseen) to the training data. Even on novel tasks that were never encountered in training, Llama-Fin remains competitive and consistently outperforms its base model across all identified capabilities.
2 Related Work
Finance LLM Capabilities Recipe Evaluation
Model Recipe Data Recipe
AdaptLLM Concept CPT CPT: Financial text + heuristic QAs constructed from the text Financial + Classification tasks + Direct answer
PIXIU Task IT IT: Financial tasks Financial + Classification tasks + Direct answer
FinLLM Concept, Task CPT → IT CPT: Financial text + Fineweb; IT: Filtered Financial tasks Financial + Classification tasks + Direct answer
FinTral Concept, Task CPT → IT → PA CPT: Financial text; IT: Financial tasks; PA: Outcome signal only Financial + Classification tasks + Direct answer
Palmyra-Fin SoTA public checkpoint, but recipe is not disclosed
Llama-Fin
Concept, IF/Chat,
Task, Reasoning
CPT+IT → PA
CPT: Financial + General text.
IT: Financial + General tasks
PA: A novel PA that leverages outcome and process signals
General + Financial tasks; Similar + Novel tasks
Classification + Open-form QA tasks
Knowledge Recall + Reasoning tasks
Direct answer + CoT
Table 1: Comparison between Llama-Fin with other finance LLMs.
Finance LLMs Table 1 summarizes popular finance-specific LLMs developed through domain-adaptive post-training. AdaptLLM (Cheng et al., 2024) focuses on CPT and constructs heuristic QA tasks from raw text, but it considers only five financial end tasks. PIXIU (Xie et al., 2023) focus on instruction-followingby creating a financial instruction-tuning dataset from diverse open financial tasks and designing a benchmark with nine end tasks for evaluation. FinLLM (Xie et al., 2024a) extends post-training across multiple stages, first performing CPT, then IT, and incorporating multi-modal capabilities via IT. It includes some general-domain data (e.g., fineWeb) but does not explore its impact systematically. Following this line, FinTral (Bhatia et al., 2024) follows this approach but is the only open FinLLM to include PA, where preference labels were given by GPT-4 on the final outcome, considering only coarse-grained signals. It also introduces multi-modality via IT and integrates tool use and retrieval in PA training. Additionally, Palmyra-Fin (Writer, 2024), a recent state-of-the-art FinLLM, reports high performance on finance tasks, particularly CFA exams [^1] https://www.cfainstitute.org/programs/cfa-program/exam, but its training recipe remains undisclosed.
Comparing to FinDaP, none of these models explicitly identify desirable capabilities as we do with FinCap, nor do they systematically explore trade-offs between CPT, IT and PA to develop a more effective training recipe. They also do not incorporate fine-grained process signals in PA to improve reasoning, as we do in FinRec. Additionally, their evaluations lack the broader range of tasks, methods, and similarities, including reasoning tasks and chain-of-thought evaluations, that we adopt in FinEval. Finally, unlike Palmyra-Fin, Llama-Fin is fully open-source, ensuring complete transparency in its training recipe, datasets, and evaluation methods, while achieving SoTA in its size category.
PA for reasoning We explore training-time approaches for improving reasoning (Jiao et al., 2024; DeepSeek-AI et al., 2025). These methods first collect trajectories and then train the LLM with the collected trajectories. This helps the model reason more accurately and faster during inference. To collect reasoning trajectories, there are two main approaches. The first is search-based (Setlur et al., 2024; Snell et al., 2024), where a trained Reward Model (RM) is used to guide a search method (e.g., Best-of-N, Beam Search) to identify the best reasoning path. The second is revision-based (Bai et al., 2022; Du et al., 2023; Madaan et al., 2023; Saunders et al., 2022), which attempts to improve the generation distribution through multi-round interactions, often by leveraging feedback from itself or another strong LLM to refine the input prompt. In practice, revision-based methods have shown mixed results and have not yet been well established as reliable for achieving improvements (Huang et al., 2024a). In contrast, search-based methods have been shown to be more effective. In FinDaP, we propose a novel training-time method that leverages a search-based trajectory collection approach, incorporating both outcome and process rewards from a Generative RM (GenRM).
3 FinDaP Framework
In FinDaP, we first identify four desired capabilities for a finance-expert LLM (FinCap, §3.1). We then develop the training recipe FinRec, which includes both the model recipe that performs CPT and IT jointly followed by PA, and the data recipe, which examines the impact of in-domain, general-domain, and mixed-domain data while introducing a novel data construction approach for PA (§3.2). We then introduce FinTrain3.3), a set of carefully curated training datasets designed to mitigate forgetting while effectively learning domain-specific knowledge. Finally, we propose an evaluation framework FinEval3.4), which considers a diverse set of tasks, ranging from familiar to novel and from general to domain-specific, while also evaluating both direct-answer and CoT methods.
3.1 Core Capabilities (FinCap)
We began by conducting a comprehensive survey of existing work and consulting two financial domain experts—a banking industry advisor and a financial industry product manager. Through this process, we identified four key fundamental capabilities essential for a finance-focused LLM: understanding domain-specific concepts to process financial language accurately, performing domain-specific tasks to solve real-world problems, reasoning effectively to analyze complex financial data, and following instructions to interact naturally in practical applications. These capabilities are deeply interconnected—reasoning depends on conceptual knowledge, while instruction-following ensures effective communication.
$\bullet$ Domain specific concepts. A domain typically includes its own specific concepts. For example, ‘bond’ in finance refers to a loan agreement between an investor and a borrower. Adapting the LLM to domain-specific concepts is crucial, as these concepts form the fundamental building blocks of domain knowledge. However, this adaptation should not come at the cost of losing knowledge about general concepts, which are essential for both domain-specific and general tasks.
$\bullet$ Domain specific tasks. While many NLP tasks, such as NER or sentiment analysis, are shared across different domains, a domain typically has its own tasks. For example, stock movement detection is primarily found in finance. Adapting LLMs to these domain-specific tasks is important, as it demonstrates how they can leverage domain-specific concepts to solve tailored tasks effectively.
$\bullet$ Reasoning. For complex tasks, reasoning with concepts is a highly desired capability in LLMs. For example, in finance, the LLM is often required to analyze a company’s financial report, involving extensive reasoning, particularly mathematical reasoning, to compute key financial concepts such as market rate or earnings per share.
$\bullet$ Instruction-Following (IF) and chat. This is a core capability for both general and domain-specific LLMs, as tasks are often presented in the form of instruction following or conversation.
3.2 FinDaP Training Recipe (FinRec)
As shown in Figure 1, FinRec consists of two recipes: the model recipe, which focuses on the training stages and losses, and the data recipe, which focuses on constructing training data.
3.2.1 Model Recipe
Previous studies often de facto treat domain-adaptive post-training as a sequential process involving, or partially involving, CPT, IT, and PA. However, our experiments with LLaMA3-8B-Inst show key trade-offs among these stages (Appendix §B). While CPT is effective at introducing domain concepts, it often leads to significant forgetting of general concepts and instruction-following capabilities. In contrast, IT strengthens instruction-following capabilities and introduces domain-specific tasks with minimal forgetting. IT alone however struggles with task generalization. PA is effective for learning reasoning but depends heavily on high-quality preference data, which can be difficult to synthesize. To address these limitations, we propose a joint CPT+IT approach, resulting in CPT+IT checkpoint. Subsequently, PA is performed with a novel trajectory collection method that provides fine-grained supervision signals.
Joint continual pre-training and instruction-tuning (CPT + IT)
In this stage, the goal is to learn domain-specific knowledge while maintaining general capabilities, such as instruction-following. It is well known that CPT can adapt the LLM to learn domain-specific concepts while IT can help learn the domain-specific and instruction-following tasks Ke et al. (2022); Wei et al. (2022). Typically CPT involves next-token prediction without masking any context tokens, and IT involves next-token prediction with instructions masked out; thus training them sequentially from an instruction-tuned LLM naturally leads to forgetting general capabilities, including instruction-following. Intuitively, if the loss function incorporates both CPT and IT, forgetting can be largely mitigated Scialom et al. (2022). [^2] This is akin to ‘replay’ method in continual learning. To achieve this, we mix CPT and IT data, effectively performing joint optimization, as the only difference between the two is whether the instruction is masked. This approach also facilitates knowledge transfer, as CPT helps the model learn domain knowledge, which can be leveraged by IT training. More importantly, since concepts learned from CPT are often inherently more generalizable due to the shared nature of concepts across tasks, jointly training CPT and IT can improve generalization without require exposure to a diverse range of tasks Wei et al. (2022), which is often impractical in certain domains, particularly long-tail ones. Since CPT datasets are typically much larger than IT datasets, we downsample CPT data to match the size of IT data, allowing for effective joint training.
Improving reasoning with preference alignment
CPT+IT improves capabilities such as in general and domain-specific concepts, tasks and IF/Chat. However, we find that the resulting model lacks in its reasoning capability, especially when it comes to complex reasoning like solving problems in CFA exams, where it is important to make each reasoning step correct. We use PA for this, which trains the model to assign higher probability mass to better generations, and has been shown to be effective in enhancing LLM reasoning capabilities (Lambert et al., 2024; Jiao et al., 2024). Specifically, we employ Direct Preference Optimization or DPO (Rafailov et al., 2023), which allows the model to learn from both positive and negative examples, providing a richer learning signal compared to SFT. We synthetically generate such data from the on-policy model, i.e., the jointly trained CPT+IT checkpoint, as it has shown the strongest performance in previous experiments. We propose a novel trajectory collection method that provides fine-grained step-level supervision signals (§3.2.2)
3.2.2 Data Recipe
While data quality and diversity are standard concerns in LLM training, we focus on two under-explored challenges: (a) the impact of in-domain, general-domain, and mixed-domain datasets on model performance at different training stages; (b) the generation of fine-grained supervision signals in PA to improve reasoning.
Mixture of in- and general-domain data
Most existing finance LLMs rely exclusively on in-domain data in post-training with the exception of FinLLM, which uses general domain data in CPT (see Table 1). Intuitively, this exclusive reliance on in-domain data can lead to forgetting of general knowledge in the original pre-trained LLM. To understand how the forgetting happens across different stages, we conduct ablation by constructing three versions of data for each training stage: in-domain, general-domain, and a mixture of both. Experiments (Appendix §B) show that the impact of forgetting decreases progressively from CPT to IT to PA, with CPT experiencing the most severe forgetting and PA the least. Guided by these findings, we adopt the mixture of in- and general-domain data for CPT+IT training to maximize both specialization and retention of essential general knowledge.
Preference data construction for reasoning
Figure 2: An overview of the proposed final answer preference (FAP) and stepwise corrective preference (SCP). In FAP, we collect trajectories from the GenRM by evaluating the entire solution (thus question marks in the intermediate steps). In SCP, we collect trajectories from the GenRM, by identifying and correcting the first erroneous step.
Existing training methods to improve reasoning primarily rely on outcome-based rewards, which provide sparse supervision and do not guide intermediate reasoning steps. At the same time, stepwise reward models can be computationally expensive if applied at every step. To strike a balance, we employ a Generative Reward Model (GenRM) and design Final Answer Preference (FAP) to efficiently collect preference signals at the final answer level (outcome reward), while also collecting Stepwise Corrective Preference (SCP) at the reasoning step level (process reward) by asking the GenRM to identify and correct the first erroneous step. By combining these two complementary strategies, our PA provides stronger supervision signals for reasoning improvements while maintaining efficiency, making it particularly suitable for domains like finance, where both accuracy and efficiency are critical. Figure 2 illustrates the proposed method.
$\bullet$ Final Answer Preference (FAP). Given a prompt and a model generated solution, we prompt the GenRM to give a holistic judgment for the entire solution using a single ‘Yes’ or ‘No’ token. We then use the correct solutions as chosen samples and the incorrect solutions as rejected samples.
$\bullet$ Stepwise Corrective Preference (SCP). Since reasoning could be complex (e.g., CFA exams) and process rewards have been shown to be more effective in such cases (Lightman et al., 2024), we further leverage the GenRM to provide step-level signals. Instead of requesting rewards at each step, which has been shown to be unnecessary in Lightman et al. (2024); Luo et al. (2024), we prompt the GenRM to identify the first erroneous step and ask it to provide a correction for that step. Using this correction, we construct a preference data sample. The input prompt is formed by concatenating the original question, the candidate reasoning steps up to the first error, and a follow-up question framed as “What is the next step?”. The chosen response of this preference sample is the newly-obtained corrected step, while the original first erroneous step is deemed as rejected response. This approach produces trajectories that focus on predicting the correct next step given a reasoning prefix, rather than requiring a prediction of the entire reasoning trajectory, as in FAP. Additional details about this approach are provided in Appendix G.
3.3 FinDaP Training Data (FinTrain)
In FinTrain, we carefully balance the trade-off between quality and quantity of the training data at each stage. Specifically, in CPT, we leverage available general-domain supervised data, such as NaturalInstrution Mishra et al. (2022) and Aqua Ling et al. (2017). Since such data has been carefully curated and cleaned for labeling, they maintain good quality. Furthermore, because of the presence of the supervised responses, they can be helpful in mitigating the forgetting issue (Scialom et al., 2022). For quantity and diversity, which is essential for learning new domain knowledge during CPT, we collect large-scale, diverse data from relevant sources, including 70 financial websites and books covering 12 financial topics, e.g., CFA exam preparation materials. We further use a strong LLM to filter out low-quality tasks based on an additive scale prompt Yuan et al. (2024). This results in a CPT dataset of approximately 8B tokens.
For IT, to promote diversity, we conduct a broad survey and source general, financial, instruction-following, and reasoning datasets from public datasets. We also include large open QA datasets like FinQA Chen et al. (2021). To ensure quality, we prioritize datasets that shown to perform well in the literature, e.g., UltraChat Ding et al. (2023). We also incorporate exercises or demonstrations from books that often contain human-written CoT. The final IT dataset consists of approximately 3M prompts.
For PA, we use CFA preparation materials as a representative source for in-domain reasoning as they cover diverse financial scenarios, em-phasize complex reasoning, and, most importantly, are derived from real-world exams. We construct preference data with FAP and SCP introduced in §3.2.2. The final PA dataset consists of approximately 32K prompts. Additional details are given in Appendix D.
3.4 FinDaP Evaluation (FinEval)
Our evaluation framework FinEval is designed to systematically assess model performance across unseen tasks. Unlike prior studies that rely on a narrow set of domain-specific tasks, FinEval categorizes tasks by similarity (similar vs. novel tasks), domain specificity (general vs. domain-specific vs. reasoning tasks), and evaluation methods (direct-answer vs. chain-of-thought). By structuring evaluations along these dimensions, In total, FinEval consists of 35 tasks and can serve as a comprehensive benchmark for the expected capabilities going beyond simple task-based evaluation. We provide details about the evaluation tasks and methods in Appendix E.
4 Experiments
Setup
Task Benchmark
Llama-Fin
8B
Llama3
Instruct
8B
Llama3.1
Instruct
8B
Palmyra
Fin
70B
Phi
3.5-mini
Instruct
3.8B
Mistral
Nemo
instruct
12B
GPT4o
Sentiment Ana. FPB (Acc) 91.13 73.09 71.55 67.11 78.04 78.25 82.16
Sentiment Ana. FiQA SA (Acc) 95.32 77.87 70.64 71.91 69.36 55.74 68.51
Monetary Policy FOMC (Acc) 64.31 56.65 54.64 63.10 58.47 57.86 67.94
Named Entity NER (Rouge1) 76.69 45.03 51.22 54.29 39.37 49.84 43.02
Abs Summ. EDTSUM (Rouge1) 53.78 11.50 12.53 21.77 19.97 12.32 18.15
Table 2: Results on similar (unseen) tasks. ‘*’ indicates that ‘GPT4o’ is used as the judge. Llama-Fin and its variant without PA (i.e., the ‘CPT+IT’ checkpoint) are highlighted in blue while the closed model is highlighted in gray. The best performing model for 8b on each benchmark is bolded. The overall best performance across all models is underlined. indicates that Llama-Fin outperforms the base Llama3-8b-inst.
Capability Domain Task Benchmark
Llama-Fin
8B
Llama3
Instruct
8B
Llama3.1
Instruct
8B
Palmyra
Fin
70B
Phi
3.5-mini
Instruct
3.8B
Mistral
Nemo
instruct
12B
GPT4o
Concept General Knowledge Recall MMLU (CoT, Acc) 47.42 48.14 47.42 54.93 45.07 49.64 63.88
AI2-ARC (CoT, Acc) 89.43 89.29 89.80 89.01 87.25 88.19 97.85
Nq-open (CoT, Acc) 19.20 18.47 22.52 19.25 6.20 17.01 27.92
Finance Knowledge Recall MMLU-Finance (Acc) 64.20 65.71 66.74 75.15 68.17 61.88 86.52
Task Finance Extractive Summ. Flare-ECTSUM (Rouge1) 34.10 35.92 35.77 33.24 35.52 37.86 35.90
ESG Issue MLESG (Acc) 40.67 36.33 36.00 39.67 38.33 32.67 45.67
Rumor Detection MA (Acc) 84.00 82.60 84.20 62.60 75.40 85.20 73.80
Stock Movement SM-Bigdata (CoT, Acc) 54.14 55.3 46.06 48.70 53.26 53.53 49.18
SM-ACL (CoT, Acc) 51.99 50.51 45.30 51.21 49.84 50.75 50.97
SM-CIKM (CoT, Acc) 54.94 55.56 48.03 52.92 50.03 53.28 49.78
Fraud Detection CRA-CCF (CoT, Mcc) 0.83 -0.32 2.73 3.12 1.20 3.94 6.16
CRA-CCFraud (CoT, Acc) 34.03 14.78 17.3 33.03 45.33 32.94 49.57
Credit Scoring Flare-German (CoT, Acc) 64.00 33.50 15.00 12.00 49.50 32.50 17.00
Flare-Astralian (CoT, Acc) 44.60 66.91 11.51 12.95 46.76 56.12 51.80
CRA-LendingClub (CoT, Acc) 68.49 52.69 25.38 23.40 48.87 21.03 65.03
Distress Ident. CRA-Polish (CoT, Mcc) 15.30 12.37 15.07 13.78 69.14 11.18 17.38
CRA-Taiwan (CoT, Acc) 40.81 12.01 35.97 52.58 69.96 57.88 8.57
Claim Analysis CRA-ProroSeguro (CoT, Acc) 35.14 96.98 44.33 56.20 25.86 32.58 96.60
CRA-TravelInsurance (CoT,Acc) 41.52 6.39 80.31 17.28 94.48 73.64 54.03
Tabular QA *Flare-TATQA (CoT, Acc) 66.61 63.43 63.70 64.21 57.70 66.40 74.90
Open QA *Finance Bench (CoT, Acc) 54.00 52.70 38.00 56.67 40.70 55.30 51.30
IF/Chat General Precise IF MT-bench (1,2 turn avg) 7.36 7.88 7.92 5.80 8.38 7.84 9.10
Reasoning Math Math Reasoning MathQA (CoT, Acc) 55.08 51.16 49.35 41.51 39.40 52.46 70.82
General Social Reasoning Social-IQA (CoT, Acc) 75.23 68.83 70.73 77.28 72.82 62.95 78.92
Common Sense Open-book-qa (CoT, Acc) 82.60 77.00 82.20 87.00 80.20 76.40 94.60
Hellaswag (CoT, Acc) 81.90 73.34 69.10 69.69 67.89 61.74 81.76
Winogrande (CoT, Acc) 70.32 62.51 66.69 74.27 72.22 65.82 85.71
PIQA (CoT, Acc) 85.85 79.82 81.45 86.72 82.05 77.91 94.34
Finance Exam CFA-Easy (CoT, Acc) 66.28 60.56 60.47 36.05 61.24 65.89 83.14
CFA-Challnge (CoT, Acc) 55.56 34.44 35.56 25.56 48.89 43.33 74.44
Table 3: Results on the novel tasks. The notations are the same as in Table 2. ‘Mcc’ refers to Matthews correlation coefficient, usually used in highly imbalanced data (Xie et al., 2024a).
We apply our method to the instruction-tuned Llama3-8b-inst, resulting in Llama-Fin. A summary of the hyper-parameters and computational resource requirements is given in Table F.1. For evaluation, we compare Llama-Fin with a wide range of baselines models, including its base model, Llama3-8b-inst, and the 8B peer, Llama3.1-8b-inst. We also include comparisons with models of other sizes, such as Phi-3.5-mini-instruct (Abdin et al., 2024) (3.8B), and Mistral-Nemo-inst (Jiang et al., 2023) (12B), as well as the closed model GPT-4o (OpenAI, 2023). Furthermore, we evaluated against the latest SoTA finance-specific LLM, Palmyra-Fin (70B) (Writer, 2024). Note that there are other financial LLMs available, such as FinMa (Xie et al., 2023), Finance-LLM (Cheng et al., 2024) and FinLLaVA (Xie et al., 2024a) that focuses on multi-modality. However, they are either not publicly available (FinLLaVA) or based on less advanced LLMs (e.g., LLaMA2). In our preliminary experiments, these models performed considerably worse than both our model and the baselines. Therefore, we have only included the SoTA financial LLM in our comparisons.
4.1 Main Results
Similar (unseen) tasks
To validate our approach, we first evaluate Llama-Fin on tasks that are similar (yet unseen) to the tasks used for training (e.g., test task EDTSUM (abstractive summarization) is similar to the training task TradeTheEvent (abstractive summarization)). From Table 2, we observe that Llama-Fin outperforms all other baselines in its size category by 10% - 25% absolute gain. It also surpasses significantly larger models, such as the finance-specific Palmyra-Fin-32K (70B). Notably, Llama-Fin also exceeds the performance of GPT-4o. These results are not very surprising since the test tasks are not entirely novel, but it demonstrates the effectiveness of our data and model recipe for domain-adaptive post-training.
Novel tasks
We now evaluate the generalization of Llama-Fin on the completely novel tasks that are also aligned to the expected capabilities (FinCap). Table 3 presents the results. Below, we summarize the key takeaways from the comparison:
$\bullet$ Llama-Fin preserves general concepts (rows 2-5). We observe that Llama-Fin performs better or remains competitive with its base model in general knowledge recall tasks, indicating that it effectively preserves general capabilities and mitigating forgetting. It performs slightly worse than the base model in finance knowledge recall (MMLU-Finance), despite our finding that the CPT benefits IT (see ablations in Appendix B.3). We hypothesize that CPT helps learn concepts that are helpful but differ from those emphasized in MMLU-Finance.
$\bullet$ Llama-Fin is effective in the majority of tasks (rows 6-22). It outperforms the base model in 13 out of 17 tasks, demonstrating that our approach can lead to models that generalize well to novel, unseen tasks requiring the same capabilities.
$\bullet$ Llama-Fin preserves IF/Chat capabilities (row 23). Llama-Fin achieves a competitive MT-Bench score compared to the base model, indicating that it effectively maintains the IF capability.
$\bullet$ Llama-Fin excels in reasoning tasks (rows 24-31). For reasoning capability, Llama-Fin significantly outperforms the base models across all considered benchmarks in a large amount (up to 20% in CFA-Challenge), indicating substantial improvements in reasoning capability.
4.2 Further Analysis and Ablations
As discussed in §3.2, we performed a number of data and model ablations in pursuit of designing the best training recipe (including parameter-efficient finetuning methods like LoRA) for Llama-Fin. Those ablations are detailed in Appendix B. In this section, we present the impact of our PA strategy in the overall post-training process.
Task Benchmark Llama-Fin
Llama-Fin
(w/o PA)
Sentiment Ana. FPB 91.13 92.99
Sentiment Ana. FiQA SA 95.32 94.47
Monetary Policy FOMC 64.31 63.10
Named Entity NER 76.69 74.33
Abs. Summ. EDTSUM 53.78 54.21
Table 4: Ablation on PA on similar (unseen) evaluation set.
Capability Domain Task Benchmark Llama-Fin 8B
Llama-Fin
(w/o PA)
Concept General Knowledge Recall MMLU 47.42 47.22
AI2-ARC 89.43 88.95
Nq-open 19.20 16.20
Finance Knowledge Recall MMLU-Finance 64.20 63.93
Task Finance Extract Summ. Flare-ECTSUM 34.10 34.41
ESG Issue MLESG 40.67 42.00
Rumor Detection MA 84.00 84.60
Stock Movement SM-Bigdata 54.14 52.04
SM-ACL 51.99 49.89
SM-CIKM 54.94 44.88
Fraud Detection CRA-CCF 0.83 0.61
CRA-CCFraud 34.03 32.32
Credit Scoring Flare-German 64.00 60.50
Flare-Astralian 44.60 51.80
CRA-LendingClub 68.49 65.96
Distress Ident. CRA-Polish 15.30 0.65
CRA-Taiwan 40.81 96.41
Claim Analysis CRA-ProroSeguro 35.14 86.57
CRA-TravelInsurance 41.52 98.50
Tabular QA *Flare-TATQA 66.61 66.43
Open QA *Finance Bench 54.00 52.00
IF/Chat General Precise IF MT-bench 7.36 7.29
Reasoning Math Math Reasoning MathQA 55.08 54.30
General Social Reasoning Social-IQA 75.23 73.64
Common Sense Open-book-qa 82.60 79.20
Hellaswag 81.90 78.92
Winogrande 70.32 67.48
PIQA 85.85 84.39
Finance Exam CFA-Easy 66.28 62.31
CFA-Challnge 55.56 35.56
Table 5: Abaltion on PA on novel evaluation set.
Table 4 presents the effectiveness of PA on similar tasks. We see that PA leads to improved performance in 3 out of 5 tasks , while not causing any significant forgetting on the other two. Considering that these tasks do not need much reasoning, this is understandable as PA primarily targets the reasoning tasks.
In Table 5, we show the same ablation for the novel tasks. In Concept (rows 9-12) and IF/chat (row 30) capabilities, removing PA often leads to worse results, indicating its effectiveness. In Task (rows 13-29), we see a mixed performance with and without PA. This is again not surprising as PA specifically focuses on reasoning tasks. Interestingly, we observe that for certain tasks (e.g., CRA-TravelInsurance, CRA-Taiwan and CRA-ProroSeguro), PA negatively impacts performance, resulting in worse outcomes compared to without PA. Even GPT-4o performs poorly in these tasks. This suggests that for some tasks, leveraging reasoning capabilities might not be beneficial, as these tasks could be inherently ‘easy’ and solvable without the need for explicit reasoning. Such observations align with prior findings (Sprague et al., 2024; Liu et al., 2024). In reasoning tasks (rows 31-38), Llama-Fin is significantly better than without PA variant, further confirming that our proposed FAP and SCP are particularly effective in improving reasoning performance beyond the already strong checkpoint of Llama-Fin (w/o PA).
5 Conclusion
We introduce FinDaP, an open SoTA finance-specific post-training framework, consists of FinCap that identifies four key capabilities, FinRec which jointly trains CPT and IT, and constructing PA preference data with stepwise signals, FinTrain that implemnts FinRec and FinEval a comprehensive evaluation setups. Under FinDaP, we develop Llama-Fin, a SoTA finance LLM. In this development, we conduct a systematic study on effectively adapting a target domain through post-training. For each stage, we reveal the distinct challenges, objectives, and effective strategies. Looking ahead, we aim to explore additional domain-specific capabilities using FinDaP.
6 Limitations
While the recipe for FinDaP and Llama-Fin are effective, the performance on novel unseen tasks still requires further improvement. For example, selectively employing reasoning capabilities only for questions that require such advanced reasoning might give better results. Additionally, the data recipe is currently based on full-scale empirical experiments, which can be time-intensive. Developing low-cost experiments to reliably indicate the effectiveness of data in post-training could streamline this process and accelerate the development iteration. It is also worth noting that the same recipe may not generalize well to other model families. Different architectures or pretraining strategies might require tailored recipe to achieve optimal results, emphasizing the need for adaptability in recipe design in future research. Finally, while we focus on the four key capabilities in finance, we acknowledge there could be additional requirements (e.g., multi-modality and sensitivity, see details in Appendix §C), and leave them for future work.
Appendix A Our Contributions
In summary, our key contributions are:
  • A comprehensive guidance for finance-specific post-training, including identification of capabilities, evaluation, data and model recipe design.
  • A systematic exploration on each stage of post-training, with an emphasis on the goals, challenges and effective approaches at each stage.
  • A novel preference alignment approach that constructs preference data using on-policy trajectories guided by outcome and process signals.
  • The above yields a new state-of-the-art finance LLM (Llama-Fin) at the 8b parameter scale.
Appendix B Ablation and Understanding FinDaP
B.1 Continual Pre-training
From Text to CPT Data. A key challenge in CPT is what kind of data we should use. Given the general and domain-specific texts introduced in §D, we can construct three versions of CPT data, CPT-In contains only the financial (in-domain) text, CPT-Gen contains only the general domain data, and CPT-Mix contains the mixture of the CPT-In and CPT-Gen.
Key Data Experiments. We conduct CPT on each of the three versions of data. As shown in Figure B.1, we observe that while CPT-In and CPT-Gen outperforms in financial (Fig B.1a) and general (Fig B.1b) tasks, respectively, CPT-Mix achieves the best overall. This is expected as CPT-In can cause catastrophic forgetting on the general tasks, while incorporating general domain concepts in CPT-Mix acts as ‘replay’ mechanism to mitigate it (Scialom et al., 2022). We can also see that none of the CPT-trained LLMs outperform their base. This is unexpected because CPT invovles post-training on more specialized data, which should enhance the performance. By analyzing the output, we attribute this issue to the model forgetting how to follow instructions effectively after CPT. To quantify this finding, we evaluate the instruction following ability of these models using MT-Bench. The two-turn average scores for CPT-Mix, CPT-In, and CPT-Gen are 1, 1, and 1.0125, respectively, while the base model, achieves a score of 7.8875. These confirm that the conventional CPT applied to a instruction-tuned LLM can cause serious forgetting on instruction-following (IF) capability. In §B.2, we will see how jointly train IT and CPT can help mitigate such forgetting issue.
Figure B.1: Average performance on selected datasets for training Llama3-8b-instruct on our CPT-In, CPT-Gen and CPT-Mix. The Y-axis represents the same performance metrics as those reported in Tables 2 and 3. The selected datasets are chosen for illustration purpose based on their ability to illustrate the general trend.
In order to expose the LLM to domain-specific concepts, we first conduct continual pre-training (CPT). In CPT, we feed plain text to the LLM and perform next token prediction.
B.2 Instruction Following
To adapt the LLM to domain-specific and IF tasks, we conduct IT. The key different between IT and CPT is that IT masks out the instruction and takes as input supervised tasks.
From Prompt to IT Data.
We introduced our prompt curation in §D. We create the responses for IT by filtering existing responses or creating new responses. For prompts with existing responses, we generally keep the original responses if they were written by a human or a strong model, such as GPT-4. We also filter out empty responses. For prompts without responses, for example, exercises extracted from books that may not have solutions provided, we generate new responses using GPT-4o. Similar to CPT data, we construct three versions of IT data, IT-In, which contains only financial (in-domain) tasks, IT-Gen, which contains only general tasks, and IT-Mix, which includes a mixture of the IT-In and IT-Gen.
Key Data Experiments. Similar to CPT, we conduct IT to each of the three versions. From Figure B.2, we observe that unlike CPT, forgetting is significantly reduced. Specifically, all versions of IT are no longer worse than their base versions, indicating that the ability to follow instructions is not as severely forgotten as in CPT. This is further supported by the MT-Bench scores, where we obtained 7.2031, 6.2094, and 7.3219 for IT-Mix, IT-In and IT-Gen, respectively, all of which are significantly better than the CPT counterparts.
Figure B.2: Average performance on selected datasets for training Llama3-8b-instruct on our IT-In, IT-Gen and IT-Mix.
We observe that IT-Mix is slightly better than other data versions, suggesting that mixing general tasks remains helpful to mitigating forgetting of general concepts and tasks, although the effect is much less pronounced compared to CPT. We also see that similar tasks improve significantly over base model while novel tasks (including financial tasks and general tasks) show little change. This indicates that, in contrast to CPT, domain has less impact in IT, but task generalization is a challenging issue.
Comparison with LoRA.
Another popular approach to adapt the LLM to specific domain is Parameter-efficient Fine-tuning (PEFT), where the LLM parameters remain fixed, and only a small set of additional parameters are trained. This approach naturally mitigates forgetting issues and is more efficient in terms of trainable parameters. However, whether it can achieve performance comparable to full-model training is unclear. In Figure B.3, we experiment with PEFT, specifically using LoRA (Hu et al., 2021), with a rank size of 128, [^3] Further decreasing or increasing the rank size did not show improvement in our preliminary experiments. For example, rank size of 32, 128 and 512 yield overall averages across 10 general tasks of 0.5267, 0.5331, and 0.5215, respectively, showing only minor differences. and compare its performance with full-model finetuning (IT-Mix). We observe that with and without LoRA performs similarly, confirming that LoRA is effective for task adaptation. However, the novel tasks still show little improvement, highlighting that task generalization still remains a significant challenge.
Figure B.3: Average performance on selected datasets for training Llama3-8b-instruct on IT-Mix with full-model finetuning (IT-Mix) and LoRA finetuning (IT-Mix (LoRA)).
A plausible reason for the lack of task generalization is that effective generalization may require exposure to a diverse range of tasks (Wei et al., 2022), which is often impractical in certain domains, particularly long-tail ones. However, concepts themselves may be inherently more generalizable due to the shared nature of concepts across tasks. Based on this, we propose adding CPT either before or concurrently with the IT stage and conduct training experiments accordingly.
B.3 Combining CPT and IT
A natural choice is to conduct CPT and IT sequentially (Lambert et al., 2024). On the one hand, this is flexible as it allows for different settings (e.g., data size) in each stage. On the other hand, it does not help prevent forgetting during the CPT stage, leaving the LLM dependent on IT to ‘recover’ its instruction-following capability. To make a more grounded decision, we conduct experiments on both sequential and joint training approaches. In joint training, an additional hyperparameter to consider is the mixture ratio. We find that down-sampling CPT data to match the size of IT data is the most effective strategy.
Figure B.4: Average performance on selected datasets for training Llama3-8b-instruct on CPT-Mix and IT-Mix jointly (CPT-Mix + IT-Mix) and sequentially (CPT-Mix $\rightarrow$ IT-Mix).
Figure B.4 illustrates the comparison between joint and sequential training. In both cases, different from IT-only results shown in Figure B.2, we see improved performance on similar and novel tasks. This supports our hypothesis that CPT can help improve the generalization of IT, as the concepts are likely able to be shared across different tasks. It is further interesting to see that even the general tasks are improved, indicating that there could be positive transfer between CPT and IT. Comparing the two, we observe that joint training outperforms sequential training across financial and general tasks, as well as similar and novel tasks, highlighting the importance of preventing forgetting of CPT and knowledge transfer between CPT and IT.
Other Sampling strategies.
Besides down-sampling, we also evaluate the performance under a ‘no-sampling’ setting. Figure B.5 shows the results. We observe that in both joint and sequential training, down-sampling yields better results on financial tasks. This is understandable because down-sampling assigns more weight to IT, which is beneficial for the financial tasks. Interestingly, we observe the opposite trend for general tasks: no-sampling performs better. We hypothesize that this is because having more CPT data helps preserve general concepts more effectively, although it may diminish instruction-following abilities.
Figure B.5: Average and selected datasets performance from down-sampling or no-sampling on CPT.
Comparison with LoRA.
In Section B.2, we showed that LoRA can effectively adapt tasks but still suffers from task generalization. While we already showed that CPT can help in full-model training setting, we now explore whether CPT can help in the PEFT setting as well. Figure B.6 presents the results of applying LoRA for IT and LoRA for both CPT and IT. Surprisingly, we find that full fine-tuning significantly outperforms the LoRA counterparts across similar and novel tasks. This finding contrasts with our previous observations in Figure B.3, where performance with and without LoRA was comparable. Our results reveal that knowledge transfer from CPT to IT, which is crucial for task generalization, requires full-model training.
Figure B.6: Average performance on selected datasets for PEFT or full model fine-tuning for CPT and IT.
B.4 Preference Alignment
Figure B.7: Average performance on selected datasets for training from ‘CPT+IT’ on PA using Trajectory-GenORM, noted as PA (GenORM) and Trajectory-GenORM + Trajectory-GenPRM, noted as PA (GenORM+GenPRM).
Figure B.8: Average performance on selected datasets for PA training from the ‘CPT+IT’ checkpoint on PA-Mix and PA-In.
Negligible Forgetting in PA. As with CPT and IT, we begin by performing an ablation study on different data versions to evaluate their effectiveness. Since the degree of forgetting diminishes from CPT to IT (as observed in §B.2), we expect it to be even less pronounced in PA. To quickly evaluate this hypothesis, we take a naive approach and create PA-Mix and PA-In by using either the provided or GPT4o generated responses (as done for IT in §B.2) as the ‘chosen’ samples and the output of ‘CPT+IT’ checkpoint as the ‘rejected’ ones, based on the prompts of IT-Mix and IT-In, respectively.
Figure B.8 shows the results after PA training for PA-In and PA-Mix from the ‘CPT+IT’ checkpoint. [^4] PA trained from Llama3-8b-instruction has shown worse results compared to training from the ‘CPT+IT’ checkpoint in our preliminary experiments, as PA requires a strong initialization checkpoint. For instance, PA-Mix from Llama3-8b-instruction achieves only 29.99 on EDTSUM, whereas the CPT+IT’ counterpart achieves 54.21. As a results, we only investigate training PA from ‘CPT+IT’ checkpoint. We observe that PA-In performs comparably to PA-Mix, indicating that it may not be essential to include general tasks to prevent forgetting of concepts or tasks, unlike the cases of CPT and IT. This suggests that PA training can focus on in-domain tasks, without requiring a broader set of general tasks or raising concerns about forgetting. Given this, we use CFA exams (Table D.2 in §D) as a representative source for in-domain reasoning because they cover diverse financial scenarios, emphasize complex reasoning, and, most importantly, are derived from real-world exams. These characteristics make them a strong proxy for a broader range of financial tasks, ensuring that the model generalizes effectively within the financial domain while simplifying the training process.
Another crucial observation is that there is not much difference even for unseen similar tasks (FiQA SA, FOMC and NER) and reasoning tasks (CFA-Easy and CFA-Challenge). This highlights the limitations of the current naive PA approach and suggests room for further improvement. In FinDaP, we propose a novel PA approach that constructs preference data guided by both outcome and process reward signals.
Appendix C Other Capabilities
Besides those core capabilities mentioned in §3.1, domains may vary significantly in their sensitivity. For instance, the medical domain is highly sensitive, requiring utmost accuracy and strict adherence to ethical considerations. In contrast, domains such as entertainment may have more relaxed requirements. Another important consideration is multi-modality, as some domains require handling multiple types of input and output formats. For example, the healthcare domain may involve processing medical images alongside textual reports, while the e-commerce domain may integrate product descriptions, images, and customer reviews into a unified response. Similarly, scientific research often combines charts, graphs, and textual analysis to present findings effectively.
Appendix D FinTrain
Capability Domain CPT Dataset Size Reference
Concept General NaturalInstrution 100,000 Mishra et al. (2022)
PromptSource 100,000 Bach et al. (2022)
Math 29,837 Amini et al. (2019b)
Aqua 97,500 Ling et al. (2017)
CREAK 10,200 Onoe et al. (2021)
ESNLI 549,367 Camburu et al. (2018)
QASC 8,130 Khot et al. (2020)
SODA 1,190,000 Kim et al. (2022)
StrategyQA 2,290 Geva et al. (2021)
UnifiedSKG 779,000 Xie et al. (2022)
GSM8K 7,470 Cobbe et al. (2021)
ApexInstr 1,470,000 Huang et al. (2024b)
DeepmindMath 379,000 Saxton et al. (2019)
DialogueStudio 1,070,000 Zhang et al. (2023)
Finance Fineweb-Fin 4,380,000 -
Book-Fin 4,500 -
Total 10,177,294
Table D.1: Summary of curated texts. New datasets released with FinDaP are color-highlighted for emphasis.
Capability Domain Task IT Dataset Size Reference
Tasks Finance Relation Cls. FingptFinred 27,600 Sharma et al. (2022)
NER FingptNERCls 13,500 Yang et al. (2023)
FingptNER 511 Alvarado et al. (2015)
Headline Cls. FingptHeadline 82,200 Sinha et al. (2020)
Sentiment Cls. SentimentCls 47,600 Yang et al. (2023)
SentimentTra 76,800 Yang et al. (2023)
Summariz. TradeTheEvent 258,000 Zhou et al. (2021)
IF/Chat General IF/Chat SelfInstruct 82,000 Wang et al. (2022)
SlimOrca 518,000 Lian et al. (2023)
UltraChat 774,000 Ding et al. (2023)
ShareGPT 100,000 Link
Finance QA FinanceInstruct 178,000 Link
FingptConvfinqa 8,890 Chen et al. (2022)
FlareFinqa 6,250 Chen et al. (2021)
FlareFiqa 17,100 Yang et al. (2023)
Reasoning Math QA OrcaMath 200,000 Mitra et al. (2024)
MetaMathQA 395000 Yu et al. (2023)
MathInstruct 262,000 Yue et al. (2023)
Code QA MagicodeInstruct 111,000 Luo et al. (2023)
Finance CFA Exam Exercise 2,950 -
Total 3,161,401
Table D.2: Summary of our curated prompts. New datasets released with FinDaP are color-highlighted for emphasis. For datasets without formal references but only a URL, we provide their links.
Continual Pre-training Text Curation
To introduce domain concepts while preserving general concepts, we curate texts for CPT. Table D.1 summarizes the texts curation datasets. Specially, for general concepts, research has shown that a ‘small’ amount of general text (as little as 1%) can effectively mitigate the forgetting issue (Scialom et al., 2022). Therefore, we focus on collecting a relatively small but high-quality set of general-domain text. To achieve this, we use verifiable text, which is text written by humans and previously used in supervised tasks in the literature. Note that this contrasts with using unverifiable web text such as C4 (Raffel et al., 2020).
For domain concept, our goal is to collect both a large volume of data and maintain high quality. Following practices from the literature on training general LLMs (Lambert et al., 2024; Gunasekar et al., 2023), we source financial texts from primarily relevant websites and books. Specifically, we source financial text from two primary resources. The first source is web text, where we filter non-financial content from the FineWeb using URLs like ‘sec.gov’ and ‘investopedia.com’. The second source is books. We manually select 10 finance-related topics (e.g., ‘economics’ and ‘management’), download books on these topics, and convert them to text using OCR (Malmgren, 2014). Since OCR can make mistakes, we further employ a strong LLM to filter out content lacking educational value or unrelated to finance. Details on the financial URLs, finance-related topics, and the prompts used for filtering is shown below:
$\bullet$ Selected Financial URLs. We curated a selection of 70 financial websites to comprehensively cover diverse aspects of finance-related content on the web. These include trusted sources from financial institutions, regulatory agencies, educational platforms, and industry-specific news outlets. This diverse collection ensures representation across sub-domains such as investment, banking, personal finance, regulatory compliance, and financial planning, offering a well-rounded foundation that can cover most of the finance content in the web.
$\bullet$ Selected Topics. We select 12 topics that are cover most of books in finance. 5 of them are from business areas, including business, Accounting, Accounting, Management, Marketing, Trading; 1 is from Mathematics, i.e., Mathematical Economics; and 4 are from Economy area, including economy, econometrics, investing, and markets. We crawled the web to downloaded the books from the corresponding topics. For CFA, we use the material provided by CFA prep providers.
$\bullet$ Prompt for Filtering the Text. We explored various prompt formats to automatically extract an financial score using an LLM and found that the additive scale by Yuan et al. (2024) worked best. Figure D.1 shows the prompt we used to filter the ‘low-quality’ text. Specifically, this prompt allows the LLM to reason about each additional point awarded, unlike the single-rating Likert scale which fits samples into predefined boxes. Then, to avoid the LLM favoring highly technical content like academia papers, we focused on financial student level knowledge. By setting a threshold of 4 (on a scale of 0 to 5) during the filtering process, we were able to also retain some high-quality financial content.
Prompt for Filtering the Text Below is an extract from a text book. Evaluate whether the book has a high financial value and could be useful in an financial setting for teaching financial students using the additive scoring system described below. Points are accumulated strictly based on the satisfaction of each criterion: - Add 1 point if the extract provides educational value for financial students whose goal is to learn financial concepts or take finacial exams. It is acceptable if quizzes are not included; however, if quizzes are present, detailed solutions and explanations must also be provided Add 1 point if the extract provides educational value for financial students whose goal is to learn financial concepts or take financial exams. It is acceptable if quizzes are not included; however, if quizzes are present, detailed solutions and explanations must also be provided. Add another point if the extract addresses certain elements pertinent to finance and aligns closely with financial standards. It might offer a superficial overview of potentially useful topics or present information in a disorganized manner and incoherent writing style. Award a third point if the extract is appropriate for financial use and introduces key concepts relevant to financial curricula. It is coherent and comprehensive. Grant a fourth point if the extract is highly relevant and beneficial for financial learning purposes for a level not higher than financial students, exhibiting a clear and consistent writing style. It offers substantial financial content, including exercises and solutions, with minimal irrelevant information, and the concepts aren’t too advanced for financial students. The content is coherent, focused, and valuable for structured learning. Bestow a fifth point if the extract is outstanding in its financial value, perfectly suited for teaching either at financial students. It follows detailed reasoning, the writing style is easy to follow and offers profound and thorough insights into the subject matter, devoid of any non-financial or complex content. The extract: <EXAMPLE>. After examining the extract, You will output a json object containing the following 2 fields: { "Justification": string // Briefly justify your total score, up to 100 words. "Score": integer // Conclude with the score }
Figure D.1: Prompt for filtering the text
Insturction Prompt Curation
Prompts represent the diverse ways users may interact with models and serves the essential component for IT and PA. Table D.2 summarizes the prompts curation datasets. Specifically, we conduct a broad survey and source general, financial, instruction-following, and reasoning tasks from public datasets. To promote diversity, we include datasets like Flare-FinQA  (Chen et al., 2021), a large open QA dataset in finance, and UltraChat (Ding et al., 2023), a dataset shown to perform well for IT in the literature (Tunstall et al., 2024; Ivison et al., 2024). Additionally, we find that exercises or demonstrations from books that were curated in §D is valuable for reasoning tasks as they usually involve challenging reasonings and come with ground truth answers and sometimes even include human-written chain-of-thought (CoT) explanations.
Figure D.2 shows the prompt we used to extract exercises from books. We carefully design the prompt to extract both the question part of an exercise, which potentially include questions, scenario and exhibits, and the answer part of the exercise, which may include answer choices and solution. In books, questions and their corresponding answers can be located far apart (e.g., the questions may appear at the beginning while the solutions are provided at the end), meaning they may not be captured within the same chunk. As a result, some questions may not have corresponding extracted answers. For such cases, GPT-4o’s generated answers are used when converting the prompt into instruction-following or preference-alignment data, as detailed in Sections B.2 and 3.2.1.
Prompt for Extracting Exercise from Book You are an educational assistant aims to extract all questions from the provided material. Look for specific indicators such as "example," "quiz," "questions," or similar terms to identify where the questions are located. If the material includes scenarios or exhibits, must include all details related to them. Do not create or derive any questions or come up with content on your own—strictly extract what is present in the material. Make sure no question is missed. If one scenario or exhibits corresponds to multiple questions, duplicate the scenarios and exhitbits so that the number of questions match the number of scenarios and exhibits. The material: <MATERIAL>. After performing these tasks, You will output a json object containing the following fields: { "Justification": "string", // A brief justification for your extractions, up to 100 words. "Questions": "string", // A list of questions extracted from the material. Only extract the exact questions presented in the text. "Scenario": "string", // A list of scenarios corresponding to the above questions. If the material does not provide the scenario place "N/A." Do not do any derivation or reference, must output the exact same, detailed and complete scenarios. The scenario may contain multiple paragraphs or even splited by the exhibits, combine them into one string. The scenario can be long, you may modify it to make it shorter, but must not change its meaning. "Exhibit": "string", // A list of exhibits or tables corresponding to the above questions. If the material does not provide the exhibit, place "N/A." Do not do any summary, or derivation or cutting, must output the exact same, detailed and complete exibits. There may be multiple exhibits involved in a scenario, combine them into one string. The exhibit can be long, you may modify it to make it shorter. Must keep the table format "Answer Choices": "string", // A list of answer choices corresponding to the above questions. If the material does not provide answer choices, place "N/A." "Answer": "string" // A list of answers corresponding to the above questions. Answers should only be included if provided in the material. If no answer is given, place "N/A." If explanations or reasoning steps or equations are included, must capture all of them. Must not answer it yourself if there is no answer provided in the material. Make sure the final number of questions equals to number of scenario equals to number of exhibits equals to number of answers }
Figure D.2: Prompt for extracting exercises from books
Appendix E FinEval
Capability Domain Task Evaluation Dataset Size Reference
Unseen - Similar
Tasks Finance Sentiment Analysis FPB 970 Malo et al. (2014)
FiQA SA 235 Maia et al. (2018)
Monetary policy Stance FOMC 496 Shah et al. (2023)
Named entity recognition NER 98 Alvarado et al. (2015)
Abstractive Summarization EDTSUM 2,000 Zhou et al. (2021)
Total 3,799
Unseen - Novel
Concept General Knowledge Recall MMLU 14,042 (Hendrycks et al., 2021)
AI2-ARC 3,548 Clark et al. (2018)
Nq-open 7,842 Kwiatkowski et al. (2019)
Finance MMLU-Finance 1,460 -
Tasks Finance Extractive Summarization Flare-ECTSUM 495 Mukherjee et al. (2022)
ESG Issue Classification MLESG 300 Chen et al. (2023a)
Rumour Detection MA 500 Yang et al. (2020)
Stock Movement Prediction SM-Bigdata 1,470 Soun et al. (2022)
SM-ACL 3,720 Xu and Cohen (2018)
SM-CIKM 1,140 Wu et al. (2018)
Fraud Detection CRA-CCF 2,280 Feng et al. (2024)
CRA-CCFraud 2,100 Feng et al. (2024)
Credit Scoring Flare-German 200 Hofmann (1994)
Flare-Astralian 139 Quinlan (1987)
CRA-LendingClub 2,690 Feng et al. (2024)
Distress Identification CRA-Polish 1,740 Feng et al. (2024)
CRA-Taiwan 1,370 Feng et al. (2024)
Claim Analysis CRA-ProroSeguro 2,380 Feng et al. (2024)
CRA-TravelInsurance 2,530 Feng et al. (2024)
Tabular QA Flare-TATQA 1,670 Zhu et al. (2021)
Open QA Finance Bench 150 Islam et al. (2023)
IF/Chat General Precise IF MT-bench 80 Zheng et al. (2023)
Reasoning Math Reasoning MathQA 2,985 Amini et al. (2019a)
General Social Reasoning Social-IQA 2,636 Welbl et al. (2017)
Common Reasoning Open-book-qa 500 Mihaylov et al. (2018)
Hellaswag 10,003 Zellers et al. (2019)
Winogrande 1,767 Sakaguchi et al. (2019)
PIQA 3,000 Bisk et al. (2020)
Finance Exam CFA-Easy 1,030 Link
CFA-Challenge 90 -
Total 91,872
Table E.1: Summary of our evaluation dataset. New datasets released with FinDaP are color-highlighted for emphasis.
With the breakdown of capabilities in §3.1, our evaluation framework consists of a suite for assessing these capabilities using development sets and unseen (held-out) evaluation sets. Our development set is directly split from the training data at each stage. Table E.1 outlines the capabilities and the evaluation benchmarks selected to cover these capabilities. Crucially, we did not examine scores on our unseen set while developing the models, which allows us to observe how much we may have overfitted to particular evaluations in our decisions around training recipe. For the unseen tasks (Table E.1), we manually review each individual dataset and have the following considerations.
$\bullet$ Benchmarking tasks. Corresponding to the capabilities, we consider a diverse set of benchmarking tasks. For concepts, we include knowledge tasks in the general domain, such as AI2-ARC (Clark et al., 2018), as well as in finance, such as MMLU-Finance (Hendrycks et al., 2021). For tasks, we consider general tasks, such as Social-IQA (Welbl et al., 2017), and domain-specific tasks, such as MLESG (Chen et al., 2023a). Notably, we intentionally include a few financial tasks such as Flare-TATQA (Zhu et al., 2021) and SM-Bigdata (Soun et al., 2022) that require understanding of tabular data, as this data format is common in this domain. For IF/Chat capabilities, we utilize popular instruction-following benchmarks, such as MT-Bench (Zheng et al., 2023). For reasoning, we include general reasoning tasks, such as MathQA (Amini et al., 2019a) and Hellaswag common sense reasoning (Zellers et al., 2019), as well as domain-specific reasoning tasks, such as CRA-ProroSeguro claim analysis (Feng et al., 2024). We also construct a new benchmark on CFA-Challenge based on CFA Level III, one of the most challenging financial exams that requires comprehensive reasoning (Khamnuansin et al., 2024; Callanan et al., 2024).
$\bullet$ Evaluation method. We split our evaluation set into two types based on their exposure to Instruction tuning (IT) data (Table E.1). The first type, Similar, includes tasks whose types have been encountered during training, even if the specific tasks themselves are unseen (e.g., a new NER task). The second type, Novel, includes tasks whose types have not been seen during training, representing entirely new challenges for the model (e.g., stock movement prediction). We use two different evaluation methods based on the nature of the benchmarks. For knowledge and NLP tasks (e.g., NER), we employ a straightforward direct answer evaluation. For reasoning tasks (e.g., CFA-Challenge), we use a 0-shot chain-of-thought (CoT) (Wei et al., 2023) answer evaluation to enhance the reliability of our evaluation. This also exposes the reasoning path, allowing us to investigate the causes of incorrect answers and enabling a more fine-grained comparison across different models.
Appendix F Summary of the Final Recipe and Hyper-parameters
Final Recipe for Llama-Fin
Continual Pre-training (CPT) and Instruction Tuning (IT)
Data 50% CPT, 50% IT
Curriculum Group 1 CPT: 50% Domain-specific Text (Web and book), 50% General text (verfiable text)
IT: 20% Domain-specific tasks, 80% General tasks
Group 2 CPT: Group 1 data + domain-specific books
IT: Group1 + Exercises extracted from books
Steps
Group 1: 3.84B tokens; Group 2: 1.66B tokens
(8,000 context length, 16 A100)
Model Intialization Llama3-8b-instruct
Attention CPT: full attention with cross-docuemnt attention masking
IT: full attention with instruction mask-out and cross-docuemnt attention masking
Optim. AdamW (weight decay = 0.1, $\beta_{1}$=0.9, $\beta_{2}$=0.95)
LR Group 1: 5e-6 with 10% warmup; Group 2: 5e-6 with 50% warmup
Batch size 128K tokens
Stop Cri. Loss of development set stops decreasing ($\approx$ 1 epoch)
Preference Alignment (PA)
Data Trajectory-GenORM and Trajectory-GenPRM
Steps 24.58 M tokens
Model Initialization CPT+IT
Loss DPO with an additional negative log-likelihood term
Optim. LR 5e-7 with 10% warmup
Batch size 32K tokens
Stop Cri. Loss of development set stops decreasing
Table F.1: Final recipe of Llama-Fin. The joint training of CPT and IT is structured into two groups, with each group undergoing joint training sequentially. The second group utilizes higher-quality data (sourced from books), following the typical curriculum training practice (Gao et al., 2024). For PA, we employ a modified DPO loss with an additional negative log-likelihood term, similar to Pang et al. (2024), as it has shown to be more effective than relying solely on the original DPO loss.
Appendix G GenRM Details
In §3.2.1, we simplified the prompt for GenRM for the purpose of illustration in Figure 2. In this section, we give full detailed of the prompt for Final Answer Preference (FAP) and Stepwise Corrective Preference (SCP) in Figure G.1 and Figure G.2, respectively.
Prompt for FAP You given a question, a reference answer and a proposed answer, you task is to determine the correctness of the proposed answer. First, extract the final answer (for example, A, B or C) from the reference answer. Second, extract the final answer from the proposed answer (for example, A, B or C). Finally, compare the two final answer to determine the correctness. Do not do any extra reasoning, must determine the correctness soley based on the given reference and proposed answer. Question: <QUESTION>
Reference Answer: <REFERENCE>
Proposed Answer: <PROPOSAL>
After performing these tasks, You will output a json object containing the following fields: { "Justification": "string", // A brief justification for your output, up to 100 words. "Correctness": "string", // If the proposed answer has the same final final answer as the reference answer (for example, both choose A or have the same answer), output correct’. Put wrong to all other cases. For example, if the proposed answer has a different final answer comparing to the reference answer, put wrong’. If the proposed answer does not explicitly give a final answer to the question, put wrong’. If the proposed answer gives more than one final answer to the question, put wrong’. }
Figure G.1: Prompt for FAP
Prompt for Prompt for SCP Given a question, a reference answer and an incorrect answer, you task is to identify the first incorrect step from the incorrect answer. The "first incorrect step" means all reasoning up to that point is accurate, but the error begins at this specific step. Question: <QUESTION>
Reference Answer: <REFERENCE>
Incorrect Answer: <INCORRECT>
After performing these tasks, You will output a json object containing the following fields: { "Justification": "string", // A brief justification for your output, up to 100 words. You need to explain (1) why the identified first incorrect step is incorrect; (2) why the reasoning up to this specific step is correct and (3) how the corrected step resolves the issue, aligning with the reference answer, maintaining the logical flow and progressing to the final answer. "First incorrect step": "string", // The explanation in the incorrect answer consists of multiple reasoning steps. Please identify the first incorrect reasoning step. It should be a piece of text directly and exactly quoted from the incorrect answer. It should be an intermediate step rather than the final answer "Reasoning up to incorrect": "string", // From the incorrect answer, give the correct reasoning steps up to the first incorrect step. This should be directly and exactly quoted from the incorrect answer. "Step correction": "string", // Replace the identified incorrect step with a single, clear, and correct step. This step should directly address and correct the error, explicitly providing the correct reasoning without requring for more information or challenging the question. It should effectively answer the question, "What is the next reasoning step?" given on the question and the identied "Reasoning up tp incorrect". It should help progress to the final answer. }
Figure G.2: Prompt for SCP